### v1

# Enhancing Memory Efficiency in Large Language Model Training Through Chronos-aware Pipeline Parallelism

Abstract.
Larger model sizes and longer sequence lengths have empowered the Large Language Model (LLM) to achieve outstanding performance across various domains. However, this progress brings significant storage capacity challenges for LLM pretraining. High Bandwidth Memory (HBM) is expensive and requires more advanced packaging technologies for capacity expansion, creating an urgent need for memory-efficient scheduling strategies. Yet, prior pipeline parallelism schedules have primarily focused on reducing bubble overhead, often neglecting memory efficiency and lacking compatibility with other memory-efficient strategies. Consequently, these methods struggle to meet the storage demands of storage capacity for next-generation LLM.
This work presents ChronosPipe, a Chronos-aware pipeline parallelism for memory-efficient LLM pretraining. The core insight of ChronosPipe is to treat HBM as a fast but small ’cache,’ optimizing and exploiting temporal locality within LLM pretraining to enhance HBM utilization. ChronosPipe introduces a pipeline scheduling strategy, Chronos-Pipe, to reduce the extrinsic overhead that disrupts the temporal locality of activations. Additionally, it leverages Chronos-Recomp and Chronos-Offload to efficiently harness the intrinsic temporal locality of activations and weights in Deep Neural Networks. Experiment results show that ChronosPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
Pipeline Parallelism, Recomputation, Offload, Memory-Efficient Schedule, Temporal Locality
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. INTRODUCTION
In recent years, Large Language Models (LLMs) have demonstrated exceptional performance across diverse domains (LLM, ), leading to an urgent demand for storage capacity during pre-training. This demand can be attributed to two main reasons outlined below: Parameter Scaling: The scaling law (ScalingLaw, ) posits that LLMs with increased parameter counts yield superior performance. This trend is exemplified by the rapid expansion from GPT-1 (0.12B parameters) (GPT1, ) to Llama 3.1 (405B parameters) (LLAMA3, ), representing a more than 3000-fold increase in model size within a mere six-year span. Sequence Length Extension: Longer sequence lengths have proven instrumental in enhancing model capabilities, particularly in multimodal applications (Multimodal, ) that necessitate the processing of extensive video and audio sequences. Consequently, the memory requirements for activations, weights, gradients, and optimizer states keep soaring.
Contemporary GPU training systems utilize high-bandwidth memory (HBM) for runtime variable storage, but it now faces significant challenges. First, HBM integration requires advanced packaging techniques such as through-silicon via (TSV) (TSV, ) and microbump (microbump, ), resulting in substantial manufacturing costs. For instance, HBM accounts for 50-60% of the total costs in NVIDIA’s H100 SXM5 module (H100, ). Moreover, next-generation HBM poses considerable problems. Increasing the count of stacking DRAM die within the restricted packaging height (720 um) necessitates hybrid bonding (HybridBonding, ), whose thermal management and yield optimization (HybridBonding_Problem, ) are extremely challenging.
Therefore, LLM training necessitates memory-efficient scheduling strategies while existing Pipeline Parallelism (PP) predominantly emphasizes performance optimization and overlooks memory constraints. Currently, most works focus on pipeline bubble reduction through fine-grained task division (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ) or utilizing tasks in other dimensions (Chimera, ; Mixpipe, ). However, these methods fail to address activation storage imbalances across pipeline stages (as shown in Fig.  1(b)) and may inadvertently increase activation and weight storage requirements. Few works notice this issue and propose offloading activations to CPUs (MegatronKwai, ) or additional GPUs (Bpipe, ). However, unlike the offloading of optimizer states, which only happens once per mini-batch, activation offloading is more frequently conducted across different micro-batches, thus raising a higher demand for offloading bandwidth.
Figure 1. (a) limited HBM capacity is difficult to meet the demands of next-generation LLM. (b) existing pipeline parallelism ignores temporal locality
Moreover, current LLM training predominantly employs hybrid parallelism, yet existing PP strategies lack optimal compatibility with other memory-efficient scheduling approaches. The ZeRO series (ZeRO, ) represents state-of-the-art data parallelism (DP) techniques, distributing model states across DP machines. In hybrid DP and PP configurations, distributing optimizer states (ZeRO-1) necessitates inter-DP rank communication only at mini-batch granularity. However, further distributing weight gradients (ZeRO-2) and weights (ZeRO-3) requires more frequent communication at micro-batch granularity, imposing higher bandwidth demands between DP ranks. Consequently, hybrid DP-PP implementations (DeepSpeed, ) typically utilize only ZeRO-1, leaving substantial potential for reducing weight and gradient storage. When designing PP strategies that are more compatible with ZeRO-2/3, such as Breadth First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), these approaches reduce DP bandwidth requirements at the cost of a significant increase in activation memory, leaving no perfect solution.
Existing PP strategies lack memory efficiency due to their failure to address an intrinsic property of Deep Neural Network (DNN): temporal locality. In these networks, forward computation precedes backpropagation, causing activations in shallower layers to be generated early but released late, thus incurring a worse temporal locality. In pipeline parallelism, where layers are distributed across machines, the first stage that contains the shallowest layer experiences peak activation storage.
Unlike previous work, we notice the issues mentioned above and introduce the concept of temporal locality to pipeline parallelism for the first time. We propose ChronosPipe, which enables the training of models 2.4 times larger than 1F1B (DAPPLE, ) on the same hardware with comparable throughput while maintaining compatibility with various memory-efficient scheduling strategies. The ChronosPipe solution comprises three components: Chronos-Pipe, Chronos-Recomp, and Chronos-Offload, which are described below.
Chronos-Pipe minimizes peak activation storage by reducing micro-batch execution time, thus expediting the backward pass that consumes activations. This, in effect, optimizes activation temporal locality. Chronos-Recomp exploits the poorer temporal locality of shallower layer activations and discards them from HBM by selectively recomputing to achieve higher efficiency. Chronos-Offload takes advantage of poorer temporal locality of deeper layer weight and discards them from HBM by offloading the process of optimizer update to CPU. Since this process overlaps with PP’s warmup and cooldown phase, it lowers demand on offload bandwidth and the processing speed of CPU.
Our main contributions are as follows.
  • We found that temporal locality in DNN is the primary cause of imbalanced activation storage in PP. This imbalance limits the size of trained models on current devices.
  • For activation memory savings, we introduce temporal locality into the PP schedule (Chronos-Pipe) and recomputation (Chronos-Recomp). Chronos-Pipe eliminates unnecessary intervals, optimizing the temporal locality of activation (Section 4.1). Chronos-Recomp discards activation with poor temporal locality from HBM by selectively recomputing shallow layers to achieve higher efficiency (Section 4.2).
  • For Model State storage capacity, we show that a ZeRO-2-compatible PP could built based on Chronos-Pipe (Section 4.3). Additionally, we designed the offload strategy based on temporal locality (Chronos-Offload), which can discard model states with poor temporal locality from HBM by offloading optimizer updates of deeper layers to the CPU (Section 5.1).
  • End-to-end evaluation on ChronosPipe is carried out on a cluster comprising up to 64 accelerators. Experiments show that Chronos-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
2. BACKGROUND AND MOTIVATION
In this section, we will introduce pipeline parallelism (PP) (Section 2.1), typical recomputation strategies (Section 2.2), and offloading strategies (Section 2.3). We will then discuss the motivation for leveraging temporal locality (Section 2.4).
To facilitate clearer explanations throughout the paper, we will use the following symbols.
$\bullet$ $m_{a}$:memory consumption for activation in whole Neural Network(exclude input embedding layer and language modeling head)
$\bullet$ $P$:# of pipeline stages
$\bullet$ $m$:# of micro-batches executed within a training iteration
$\bullet$ $T_{fwd}$: execution time of forward pass in one micro-batch
$\bullet$ $T_{bwd}$: execution time of backward pass in one micro-batch
2.1. Pipeline parallelism
Figure 2. (a) Interleave-1F1B schedule. (b) ZeRO-compatible pipeline schedule
Pipeline Parallelism (PP) divides the model into multi-layer blocks mapped across different PP stages. These stages work together to train all samples within a mini-batch, enabling mini-batch Stochastic Gradient Descent. During the forward pass, samples in each mini-batch begin computation at the first stage, with subsequent blocks executing their computations in a pipelined manner until the final forward computation completes at the last stage. In the backward pass, gradients for each mini-batch flow in the reverse direction. As a result, PP only requires a single tensor to be sent and received per block, significantly reducing communication bandwidth demands. This approach is often deployed across nodes and is a key component in hybrid parallelism.
Pipeline parallelism (PP) at the mini-batch level suffers from significant pipeline bubbles, prompting the common practice of splitting mini-batches into smaller micro-batches. By dividing the workload, the first stage, after processing the initial micro-batch, can immediately start on subsequent micro-batches rather than idling, thus reducing pipeline bubbles (GPipe, ). When the number of micro-batches increases, the classic 1F1B (one-forward-one-backward) scheduling strategy (DAPPLE, ) is often used, as shown in Fig.  1(b). During the steady phase, a forward computation block is initiated only after a backward block finishes, avoiding the simultaneous launch of multiple forward passes in the warm-up phase, which would otherwise increase peak memory usage. However, this approach creates an imbalance in activation memory usage across PP stages. In a single micro-batch, the first stage starts forward computation early but is the last to complete backward passes and release memory. As a result, the first stage reaches peak activation memory of $m_{a}$, while the last stage’s activation memory usage is only $\frac{m_{a}}{p}$. Additionally, Fig.  1(b) illustrates that tasks with dependencies between adjacent PP stages execute sequentially, leading to non-overlapping point-to-point (P2P) communication. Some scheduling methods address this by inserting independent task blocks from other micro-batches between dependent blocks, enabling overlapping P2P communication. This delay caused by dependency constraints is called the ”interval” delay in this paper.
Blocks can be broken down into smaller tasks to reduce pipeline bubbles further, though this may increase memory requirements. The Interleaved-1F1B approach (Interleave-1F1B, ) divides blocks into more chunks and schedules multiple chunks launches during the warmup phase, as shown in Fig. 2(a). When the workload is divided into v chunks (assigning layer 1, $p+1$, …, $(v-1)p+1$ to stage 0; layer 2, $p+2$, …, $(v-1)p+2$ to stage 1, and so forth), the bubble size is reduced to $\frac{1}{v}$ of that in the traditional 1F1B, but peak activation memory increases to $m_{a}(1+\frac{p-1}{pv})$. The Zero Bubble approach (ZeroBubble, ) goes a step further by splitting the backward pass (BP) into two phases: activation gradient computation (BPA) and weight gradient computation (BPW). This arrangement theoretically enables a bubble-free pipeline. However, the activation gradients generated by BPA are required to be retained until BPW is completed, which raises memory demands. Additionally, splitting the BP disrupts the original reuse of activation gradients for all-gather operations between BPA and BPW, and reduces the overlap between computation and communication.
Designing a ZeRO-compatible pipeline schedule is challenging due to the trade-off between DP communication bandwidth and activation memory storage. Communication between DP ranks at the micro-batch level is required for a schedule to be compatible with ZeRO-2 or ZeRO-3. To overlap this communication, existing methods, such as Breadth-First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), use grouped execution of micro-batches, as shown in Fig. 2(b). This approach processes multiple micro-batches simultaneously, allowing communications of weights and weight gradients can be overlapped with computations, but it also causes a substantial increase in activation memory usage. Since weights and weight gradients are stored in 16-bit and 32-bit precision, respectively, achieving ZeRO-3 PP requires at least a 1.5x increase in DP communication bandwidth or an increase in the number of micro-batches per group compared to ZeRO-2 PP. Consequently, some implementations choose to use only ZeRO-2 PP (LLAMA3, ).
2.2. Recomputation
Recomputation allows only a subset of activations generated in the forward pass (called checkpoints) to be retained, with all other activations regenerated during the backward pass, trading off memory usage for additional computation. Previous work (ChainTrainingMem, ) has explored optimal checkpoint placement, given a fixed number of checkpoints, to minimize activation memory, achieving memory savings at sub-linear cost for chain structure of DNN. Additionally, the benefits of recomputation vary across different operators. Recent studies suggest recomputing only the attention part in transformers (SeqParallel, ) or recomputing some non-linear operators (MegatronKwai, ) to reduce memory usage with minimal computational overhead effectively. So far, it is still expensive to recompute the projection operator in LLM, which dominates the remaining activation memory usage. Moreover,  (DTR, ) further introduced temporal locality-aware recomputation principle and demonstrated its efficacy in data parallelism. Although pipeline stages inherently exhibit temporal locality variations in PP, coarse-grained PP fails to leverage intra-block temporal locality variance for recomputation optimization within the simplistic structure of LLMs. Existing methods thus fail in single-chunk PP. Moreover, Direct application of this principle to multi-chunk PP (e.g., interleaved 1F1B scheduling) risks inducing steady-phase bubbles, resulting in significant throughput degradation.Therefore,  (DTR, ) fail to solve imbalanced memory in PP.
2.3. Offloading Strategy
Offloading techniques allow tensors to be moved from memory-constrained GPUs to other GPUs or CPUs, trading memory usage for data movement. In BPipe (Bpipe, ), GPU offloading is used to address activation memory imbalances in pipeline parallelism. However, this approach requires high-bandwidth intra-node communication links like NVLink rather than slower inter-node links like Ethernet or InfiniBand. CPU offloading utilizes PCIe to offload activation or model states. Theoretically, CPU offloading for activation memory reduction is well-suited for long-sequence training (MegatronKwai, ). However, limited PCIe bandwidth may leave some activation in HBM in practice. For model states, previous work, such as ZeRO-Offload (ZeRO-Offload, ), offloads the process of optimizer-step to the CPU, storing both optimizer states and master weights in CPU-side DRAM. To minimize GPU idle time caused by latency in the CPU’s optimizer step, this approach imposes significant requirements on both offload bandwidth and CPU computational capacity. Overall, Recent studies still have high bandwidth requirements for the Offloading Strategy.
2.4. Motivation: Temporal-Locality Matters
Figure 3. forward and backward pass of deep neural network
In the training of DNN, the forward pass is conducted first, followed by backpropagation, which naturally contains differences in temporal locality. As shown in Fig. 3, activations of shallow layers are generated first during the forward pass but are the last to be released in the backward pass, resulting in poor temporal locality. In contrast, weights of deep layers can be updated early during backpropagation, though they are not needed until the end of the forward pass, providing ample time for updating the optimizer state—a property that can be further leveraged.
In pipeline parallelism, temporal locality reveals some additional interesting phenomena:
Higher Peak Activation Storage Requirements: The PP stage responsible for shallow-layer activations releases them last, which increases the activation storage requirements. Additionally, intervals can exist between blocks with dependencies, and these intervals can accumulate across multiple blocks, further worsening the temporal locality for shallow layers. This is the reason that the interleaved 1F1B schedule results in a higher peak storage demand.
Skewed Distribution of Peak Activation Storage: As shown in Fig.  2(a), the peak activation storage in the interleaved 1F1B schedule tends to occur at Stage 0. When the chunk size is set to 2, the ratio of peak activation storage between shallow and deep layers is $(2p-1):p$, indicating an obvious bias.
Sufficient Time for Process of optimizer-step: Beyond the natural advantage of chain-structured networks providing additional time for optimizer update on deep-layer weights, increasing the PP will provide more sufficient time for optimizer-step. Large PP does not impact the execution latency of each micro-batch while the amount of model state that needs updating per stage is reduced, thereby lowering the communication and processing demands for weight updates.
3. OVERVIEW OF ChronosPipe
Figure 4. Overall scheduling of ChronosPipe. The core idea behind ChronosPipe is to treat HBM as a fast but limited ”cache.” Chronos-Pipe enhances activation temporal locality in PP scheduling while Chronos-Recomp and Chronos-Offload discard activation and weight with poor temporal locality from HBM.
Inspired by the principle of efficient storage utilization in Cache, ChronosPipe applies cache’s core concepts to HBM. The central idea of ChronosPipe is to optimize and leverage temporal locality during LLM pretraining, enabling more efficient utilization of the HBM.
ChronosPipe classifies the factors affecting temporal locality in LLM training into two main types. First, temporal locality varies across different layers within the same network; this influence is referred to as intrinsic temporal overhead, capturing the impact on temporal locality inherently introduced by the model’s structure. Second, the same layer may display different temporal locality depending on the scheduling strategy used; this influence is termed extrinsic temporal overhead, capturing the effects on temporal locality introduced by scheduling differences.
To address extrinsic temporal overhead, ChronosPipe draws on the first critical insight from Cache: optimize locality in applications. As illustrated in Fig.  4, ChronosPipe introduces a PP scheduling strategy called Chronos-Pipe to advance the backward passes of both shallow and deep layers, minimizing the lifespan of activation. Detailed explanations of Chronos-Pipe are provided in Section 4.1. This scheduling strategy reflects ChronosPipe’s approach to optimizing temporal locality in LLM training.
While scheduling can only reduce extrinsic temporal overhead, addressing intrinsic temporal overhead still requires additional strategies. To tackle this problem, ChronosPipe draws on a second critical insight from Cache: discard data with poor locality from Cache. As illustrated in Fig.  4, ChronosPipe removes data with poor intrinsic temporal locality from HBM. This involves Chronos-Recomp, a recomputation strategy for shallow-layer activations, and Chronos-Offload, an offloading strategy for deep-layer weights. Detailed explanations of Chronos-Recomp and Chronos-Offload are provided in Sections 4.2 and 5.1, respectively. Through recomputation and offloading, ChronosPipe leverages temporal locality to improve memory efficiency in LLM training.
4. Chronos-aware SCHEDULE FOR ACTIVATION MEMORY SAVINGS
4.1. Chronos-Pipe: Chronos-aware Pipeline
Figure 5. (a) Chronos-Pipe Pipeline Schedule(chunk size=2). (b) Analysis of overhead of P2P communication in Chronos-Pipe.
Chronos-Pipe Schedule
The primary goal of Chronos-Pipe is to mitigate the impact of interval accumulation in pipeline scheduling, which worsens activation temporal locality. To leverage the temporal locality differences across layers, we further split blocks into multiple chunks and prioritize completing backpropagation as early as possible. Unlike interleaved 1F1B, we aim to execute tasks with dependencies in adjacent stages as tightly as possible to minimize the negative effects of interval accumulation across multiple tasks on shallow-layer temporal locality. As shown in Fig.  5(a), with a chunk size of 2, Chronos-Pipe’s scheduling introduces an interval only between the two chunks in both the forward and backward passes.
Theoretical Analysis
A further theoretical analysis of Chronos-Pipe’s impact on activation storage requirements and P2P communication is provided. For simplicity, we assume $T_{bwd}=2T_{fwd}$ and set $T_{c}$ for a single P2P communication to zero unless stated otherwise. With these assumptions, basic unit of time $T_{unit}$ in the pipeline scheduling diagram (e.g., Fig.  5(a)) is $\frac{T_{fwd}}{2p}$, and the activation storage for each task is $\frac{m_{a}}{2p}$.
In Chronos-Pipe, activation storage can be estimated by analyzing the activation lifespan, approximately 75% of $m_{a}$. As shown in Fig.  5(a), Stage 0 launches chunk 2 of the fourth micro-batch before releasing the activation storage from chunk 2 of the third micro-batch. Since tasks on a single stage execute in cycles of $6T_{unit}$, the accumulated activation storage for chunk 2 can be represented as $\lceil\frac{T_{life\_chunk2}}{6T_{unit}}\rceil$, where $T_{life\_chunk2}$ denotes the lifespan of chunk 2’s activations, precisely the time from the completion of its forward pass to the beginning of the corresponding backward pass at PP stage 0. Similarly, the accumulation of activation storage for chunk 1 can be expressed as $\lceil\frac{T_{life\_chunk1}}{6T_{unit}}\rceil$. Additionally, we observe that the forward execution times for both chunks $T_{fwd\_chunk1}$, $T_{fwd\_chunk2}$ are each $pT_{unit}$, while the backward execution times $T_{bwd\_chunk1}$, $T_{bwd\_chunk2}$ are each $2pT_{unit}$. There is also an extra interval $T_{fwd\_interval}$ between the forward chunks, derived from steps 1, 2, and 3 in Fig.  5(a) as $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$. Similarly, an additional interval $T_{bwd\_interval}$ exists between the backward chunks, derived from steps 4, 5, and 6 in Fig.  5(a) as $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$. Thus, for Stage 0, the lifespans of the activations for chunk 1 and chunk 2, denoted $T_{life\_chunk1}$,$T_{life\_chunk2}$ can be expressed as:
(1) $T_{life\_chunk2}=T_{fwd\_chunk2}+T_{bwd\_chunk2}-2T_{unit}$
(2) $T_{life\_chunk1}=$ $T_{fwd\_chunk1}+T_{fwd\_interval}+$
$T_{life\_chunk2}+T_{bwd\_interval}+T_{bwd\_chunk1}$
As a result, the peak activation storage requirements for chunks 1 and 2 are $\lceil\frac{2}{3}+\lceil\frac{(p-3)}{6}\rceil+\lceil\frac{(2p-3)}{6}\rceil+\frac{p}{2}\rceil\frac{m_{a}}{2p}$,$\lceil\frac{(3p-2)}{6}\rceil\frac{m_{a}}{2p}$ respectively. When P is large, these values approach ${}\frac{m_{a}}{2},\frac{m_{a}}{4}$, making the total peak activation storage approximately 75% of $m_{a}$.
In Chronos-Pipe, when the task is divided into v chunks, the bubble overhead from P2P communication is approximately v times that of the 1F1B schedule. The scheduling of Chronos-Pipe with two chunks under the condition $T_{c}\neq 0$ is considered to illustrate this. As shown in Fig.  5(b), the Chronos-Pipe execution time can be divided into three phases: $T_{warmup},T_{steady}$ and $T_{cooldown}$.
$\bullet$ The warmup phase primarily includes the forward pass of the first micro-batch, calculated as $(2p+2)T_{unit}+(2p+1)T_{c}$.
$\bullet$ The Steady phase covers the forward pass of chunk 1 for the remaining $(m-2)$ microbatches, totaling $(m-2)$$(6T_{unit}+8T_{c})$.
$\bullet$ The cooldown phase mainly involves the backward pass of the last micro-batch, calculated as $(4p+4)T_{unit}+(2p+3)T_{c}$.
Adding the time of these three phases together, the total time is $6(m+p-1)T_{unit}+(4p+8m-12)T_{c}$. Exactly, Fig.  5(b) shows that the warmup and cooldown phases can each remove one $T_{c}$ of bubble overhead, leading to an adjusted total time of $6(m+p-1)T_{unit}+(4p+8(m-2)+2)T_{c}$. In comparison, the execution time for 1F1B is derived as $6(m+p-1)T_{unit}+(2p+4(m-2))T_{c}$. From this, we can see that Chronos-Pipe with two chunks has a P2P communication overhead slightly greater than twice that of 1F1B. This is because P2P communication between Stage 0 and Stage (P-1) appears in the critical path of the Chronos-Pipe schedule during the warmup and cooldown phases. This analysis holds for other values of $v$, where the P2P communication bubble overhead will slightly exceed v times that of 1F1B. Further analysis shows that the theoretical bubble overhead for Chronos-Pipe with two chunks is:$\frac{6(p-1)T_{unit}+[4p+8(m-2)+2]T_{c}}{6(p-1+m)T_{unit}+[4p+8(m-2)+2]T_{c}}$ whereas the bubble overhead for 1F1B is: $\frac{6(p-1)T_{unit}+[2p+4(m-2)]T_{c}}{6(p-1+m)T_{unit}+[2p+4(m-2)]T_{c}}$Set $T_{c}=0$, the bubble overhead for Chronos-Pipe matches that of 1F1B. When $T_{c}=0.05T_{unit},m=128,p=4$, the theoretical bubble overheads are 8.27% for Chronos-Pipe and 5.37% for 1F1B. Due to the increased P2P communication in Chronos-Pipe, the Model Flop Utilization (MFU) is slightly reduced.
Implementation of Chronos-Pipe
In Chronos-Pipe’s implementation, partial asynchronous communication is introduced to reduce bubble size. In mainstream frameworks like DeepSpeed and Megatron, the 1F1B configuration uses synchronous P2P communication. However, in Chronos-Pipe, there can be significant intervals between chunks during forward and backward passes. If synchronous P2P communication is used in these intervals, much of the time would be spent waiting for communication to complete. To address this, Chronos-Pipe modifies P2P communication between chunks to be asynchronous, thereby achieving higher throughput. As illustrated in Fig.  5(b), after Stage 0 completes the backward computation for chunk 2, it can proceed with subsequent tasks until just before Stage (P - 1) begins the corresponding backward pass for chunk 1, at which point the P2P communication is completed. Other P2P communications remain synchronous, which is consistent with mainstream frameworks.
Figure 6. (a) build Chronos-Recomp based on Chronos-Pipe. (b) Chronos-Recomp Schedule with PP stage is set to 8.
4.2. Chronos-Recomp: Chronos-aware Recomputation
Recompute Shallow Layer First!
Chronos-Pipe aims to shorten the lifespan of activations in shallow layers by reducing the interval. Yet, it cannot fully eliminate the intrinsic differences in temporal locality across layers within DNN. We can further mitigate this issue by discarding data with low temporal locality from the fast but limited HBM. This approach suggests prioritizing the recomputation of shallower layers—an approach we refer to as Chronos-Recomp—when recomputation is considered.
Chronos-Recomp can be seamlessly built based on Chronos-Pipe. We can reserve time for recomputation just before the backward computation block of the shallow layers, as shown in Fig.  6(a). This approach introduces regular intervals during the forward pass without changing the overall pipeline scheduling. Therefore, Chronos-Recomp can also support partial recomputation by controlling the length of the interval. As discussed in the theoretical analysis of Chronos-Pipe, the intervals between Chronos-Pipe chunks depend on the value of P. As P increases, dependency conflicts may arise in this approach. However, these conflicts can be avoided by delaying the forward computation launch of chunk 2, as illustrated in Fig.  6(b) for Chronos-Recomp scheduling with P=8. This adjustment does not affect the bubble ratio and memory usage, which will be explained in the following theoretical analysis of Chronos-Recomp.
Theoretical Analysis
We analyze activation storage requirements using full recomputation of shallow layers as an example. With a chunk size of 2, the remaining activation storage is capped at only 25% of $m_{a}$. Observing micro-batch 2 in Fig.  6(b), the total intervals introduced during the forward execution of chunk 2 is $\lceil\frac{(p-1)}{2}\rceil T_{unit}$. Therefore, for Stage 0, the lifespan of chunk 2’s activations is given by:
(3) $T_{life\_chunk2}=(3p+\lceil\frac{(p-1)}{2}\rceil-2)T_{unit}$
Here, $(3p-2)T_{unit}$ represents the ideal time from the start of chunk 2’s forward computation to the arrival of the backward computation, assuming no interval is present. Moreover, since tasks in pipeline parallelism repeat with a period of $7T_{unit}$, the required number of activation storage blocks for chunk 2 is $\lceil\frac{T_{life\_chunk2}}{7T_{unit}}\rceil$, which simplifies to $\lfloor\frac{p}{2}\rfloor$. Thus, with Chronos-Recomp’s full recomputation for chunk 1, the remaining activation storage is only 25% of $m_{a}$. By comparison, standard recomputation in 1F1B with the same computation budget requires 50% of $m_{a}$ for storage, making Chronos-Recomp 1.5 times more storage-efficient than standard recomputation.
Under these conditions, we can also evaluate the total execution time, where Chronos-Recomp is similar to 1F1B with recomputation, as shown in Fig.  6(b). The warm-up phase, $T_{warmup}$, represents the time required to complete the forward computation of chunk 1, taking $2pT_{unit}$. The steady phase, $T_{steady}$, involves the forward computation of chunk 2 for the remaining $(m-1)$ microbatches, consuming $7(m-1)T_{unit}$. During the cool-down phase, the backward computation of chunk 1 for the final micro-batch requires $4pT_{unit}$. This results in a total execution time of $[6p+7(m-1)]T_{unit}$. By comparison, the execution time for 1F1B with 50% recomputation is $7(m-1+p)T_{unit}$. Notably, Chronos-Recomp’s forward computation allows for overlap with half of the P2P communications. In practical scenarios, where $T_{bwd}\leq 2T_{fwd}$, the P2P communication of gradient in the backward pass of chunk 2 can overlap with recomputation. Consequently, the impact of P2P communication on Chronos-Recomp and 1F1B is similar.
As P increases, dependency conflicts can arise within the Chronos-Recomp scheduling scheme. We’ve analyzed this situation in detail. When P increases (e.g., $P\geq 8$), delaying the launch of the forward computation for chunk 2 effectively prevents dependency conflicts between the forward computations of the two chunks. Additionally, we demonstrate in Appendix A that for $P\leq 40$, delaying the computation of chunk 2 by just one round is sufficient to avoid forward dependency conflicts between chunks—a parallelism level that significantly exceeds typical configurations in mainstream settings. Notably, chunk 2’s forward execution begins later than chunk 1’s, while its backward pass completes earlier. This scheduling adjustment, therefore, does not impact the critical path and does not affect the overall execution time. Moreover, the peak activation storage analysis we conducted does not impose any limitations on the value of P, making the conclusions widely applicable across various P values.
4.3. Compatible with ZeRo-2 Based Pipeline Schedule
Figure 7. build Zero-2-compatible schedule based on Chronos-Pipe
Chronos-Pipe can be further adapted to support a ZeRO-2-compatible PP. As illustrated in Fig. 7, the modified Chronos-Pipe repeatedly launches the workload for the current chunk’s forward computation. This increases the overlap time available for ZeRO-2’s communication between DP ranks. Additionally, this adjustment has minimal impact on activation storage, enabling the launch of additional workloads of the same chunk to reduce communication demands between DP ranks. Furthermore, When recomputation needs to be considered, such as long sequence training, Chronos-Recomp can offer higher recomputation efficiency for this ZeRO-2-compatible PP scheduling approach.
5. Chronos-aware SCHEDULE FOR MODEL STATE MEMORY SAVINGS
Temporal locality can be applied not only to activation memory usage optimization but also to model state store optimization.
5.1. Chronos-Offload: Chronos-aware Offload
Chronos-Offload Scheduling Principle
The central idea behind Chronos-Offload is to offload model states with poor temporal locality to the CPU. In DNN, weights of deeper layers complete gradient computation first, allowing them to be updated earlier. However, these weights are the last to be used in the forward pass, giving them the poorest temporal locality. In the two-chunk Chronos-Pipe design, this is reflected by chunk 2 starting its forward computation later than chunk 1 but finishing its backward pass earlier. As a result, chunk 2’s weight updates can be offloaded to the CPU, while chunk 1’s weight updates remain on the GPU. This approach allows us to overlap chunk 2’s weight update process with the idle time between the end of its backward pass in the current mini-batch and the start of its forward pass in the next mini-batch.
This approach aligns well with the Optimizer Step phase’s specific computation and storage characteristics. In this phase, gradients are used to update first- and second-order momenta and the master weight (32-bit). The updated master weight is then quantized to produce a 16-bit quantized weight. Since this process involves only element-wise operations, it doesn’t leverage the high computational power of Tensor Cores, resulting in inefficient GPU utilization. On the other hand, server CPUs, with their SIMD processing capabilities, are well-suited to the computational demands of the Optimizer Step. By offloading weight updates to the CPU, the first- and second-order momenta and the master weight can be kept in CPU-side DRAM, with only the quantized weight and gradients stored on the GPU. This reduces the GPU’s storage requirements to roughly one-third of the total Model State. Through Chronos-Offload, we introduce temporal locality into the Model State Offload strategy and make it more suitable with PP, enabling complementary use of the CPU and GPU’s respective computing and memory strengths while minimizing impact on GPU computational efficiency.
Implement of Chronos-Offload
Chronos-Pipe offers seamless compatibility with Chronos-Offload in pipeline parallelism compared to other methods. In existing approaches, the cooldown phase immediately transitions from chunk 2’s backward pass to chunk 1’s backward pass, eliminating scheduling bubbles. A similar issue arises in the warm-up phase, as seen in the interleaved 1F1B structure (Fig.  2(a)). Supporting Chronos-Offload in this setup requires modifying the PyTorch backend to asynchronously manage chunk 2’s weight updates on the CPU, including gradient offloading, optimizer updates, and updated weight uploads. These intrusive changes increase programming complexity and risk resource conflicts, reducing compatibility with other scheduling strategies. In contrast, Chronos-Pipe naturally supports Chronos-Offload by introducing bubbles between tasks. During the cooldown phase, bubbles exist between the end of chunk 2’s backward pass and the start of chunk 1’s, with a similar effect in the warm-up phase. These bubbles provide ideal opportunities for Chronos-Offload without interference. This behavior stems from Chronos-Pipe’s optimization of temporal activation locality, minimizing intervals between pipeline tasks while naturally creating sufficient bubbles during warm-up and cooldown phases.
Figure 8. seamless compatibility with Chronos-Offload in Chronos-Pipe
Chronos-Pipe utilizes these bubbles to enable smooth integration with Chronos-Offload, as shown in Fig. 8. During the cooldown phase, gradients of chunk 2 are offloaded, and the corresponding optimizer updates are performed in the bubbles between chunk 2’s backward pass and chunk 1’s. Similarly, the upload of quantized weights is scheduled during bubbles between forward passes of the two chunks in the next minibatch’s warm-up phase. While this approach doesn’t fully optimize Chronos-Offload’s efficiency, we demonstrate that its effectiveness improves with higher levels of pipeline parallelism (PP), data parallelism (DP), and longer sequence lengths. Furthermore, the optimizer step involves synchronized operations, such as gradient norm calculations, which aggregate all gradients. Since Chronos-Offload splits optimizer updates between the CPU and GPU, synchronization overhead must be addressed. To mitigate this, a post-optimizer update validation strategy, similar to (ZeroBubble, ), can be employed to manage these challenges effectively.
Scalability with Increasing DP, PP, and Sequence Length
Two conditions must be met to complete the Chronos-Offload process within the available ’bubbles.’ These conditions are more easily satisfied as the degree of PP and sequence length increase. The first condition requires the gradient offload and the CPU optimizer updates to be completed within the cooldown phase’s ’bubbles.’ In Chronos-Pipe, the interval between backward computations of chunks is $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$, which results in a delay of $\lceil\frac{(2p-3)}{6}\rceil$ rounds for chunk 1 relative to chunk 2. This aligns the backward computation of the first micro-batch for chunk 1 with the (1+$\lceil\frac{(2p-3)}{6}\rceil$) th one of chunk 2. Thus, we define:
(4) $T_{avaliable\_offload}=(p-\lceil\frac{(2p-3)}{6}\rceil-1)\frac{T_{bwd}}{2p}$
Here, $T_{avaliable\_offload}$ represents the idle time available between the end of the backward computation for chunk 2 and the start of backward computation for chunk 1 during the cooldown phase. Ideally, this time would be sufficient to complete the gradient offload and the CPU update calculations. It follows that:
(5) $\frac{T_{step}}{2p}\leq T_{avaliable\_offload}\Rightarrow\frac{T_{step}}{T_{bwd}}\frac{1}{(p-\lceil\frac{(2p-3)}{6}\rceil-1)}\leq 1$
Where $T_{step}$ denotes the total time required for offloading gradients in all layers of DNN and performing CPU optimizer updates. As $P$ increases, the impact of the constant terms decreases, simplifying the equation to $\frac{T_{step}}{T_{bwd}}\frac{3}{2p}\leq 1$ for large $P$, nearly offering a linear increase in idle time, thus facilitating easier fulfillment of this condition. The ratio $\frac{T_{step}}{T_{bwd}}$ is constant for a given model and is influenced by factors such as sequence length, batch size, offload communication bandwidth, and CPU processing speed. Longer sequences and larger batch sizes further decrease $\frac{T_{step}}{T_{bwd}}$, linearly increasing the idle time available for Chronos-Offload, even without considering the attention operator.
The second requirement is that quantized weight uploads must be completed within the ’bubbles’ of the warm-up phase. Utilizing a similar analytical approach, we observe that the forward pass of chunk 2 is delayed by $\lceil\frac{(p-3)}{6}\rceil$ rounds relative to chunk 1. Consequently, we derive:
(6) $T_{avaliable\_upload}=(p-\lceil\frac{(p-3)}{6}\rceil-1)\frac{T_{fwd}}{2p}$
(7) $\frac{T_{upload}}{2p}\leq T_{avaliable\_upload}\Rightarrow\frac{T_{upload}}{T_{fwd}}\frac{1}{(p-\lceil\frac{(p-3)}{6}\rceil-1)}\leq 1$
Here, $T_{avaliable\_upload}$ represents the idle time during the warm-up phase available for uploading quantized weights, and $T_{upload}$ denotes the total time required to upload quantized weights of all layers. As $P$ increases significantly, this relationship simplifies to $\frac{T_{upload}}{T_{fwd}}\frac{6}{5p}\leq 1$. Similar to the first condition, this requirement becomes easier to meet with increased pipeline stages p and longer sequence lengths. Thus, Chronos-Offload exhibits strong scalability with increasing degrees of pipeline parallelism (PP) and sequence lengths, which is suitable for future development trends.
Additionally, Chronos-Offload scales effectively with increased Data Parallelism (DP). As the level of DP increases, the amount of weights each DP rank needs to update decreases, which in turn reduces the time required for gradient offloading and optimizer computations.
5.2. Compatible with Other Schedule
Chronos-Offload is designed to be compatible with a variety of scheduling strategies. Our approach exclusively leverages the ’bubbles’ created within Chronos-Pipe, eliminating competition over communication bandwidth and CPU with other scheduling approaches. As a result, it seamlessly supports many parallelism strategies, including tensor parallelism, context parallelism, and data parallelism. This flexibility makes it widely applicable across diverse distributed parallelism training frameworks.
Chronos-Offload is also fully compatible with the CPU Activation Offloading scheme. It performs Model State Offloading in the interval between the completion of chunk 2’s backward computation and the onset of the forward computation for the next mini-batch of chunk 2. This process exclusively utilizes the PCIe upload bandwidth during the warm-up phase and the PCIe offload bandwidth during the cool-down phase. This approach does not interfere with CPU Activation Offloading; rather, it enhances the efficient utilization of the PCIe bidirectional bandwidth.
Therefore, Chronos-Offloading effectively utilizes the temporal locality of weights while also maximizing the use of the prevalent ’bubbles’ during pipeline parallelism’s warm-up and cool-down phases. Consequently, it can seamlessly integrate with numerous scheduling optimization strategies.
6. EXPERIMENTAL RESULTS
In this section, we seek to answer the following questions:
$\bullet$ How well does ChronosPipe perform in memory and throughput?How does it compare with other recomputation-based solutions?
$\bullet$ How does ChronosPipe perform across varying sequence lengths and model architectures (input embedding, attention, activation function)?
$\bullet$
What is the actual impact of P2P communication on ChronosPipe? Can the scalability of Chronos-Offload truly maintain performance without degradation?
$\bullet$ What potential benefits could be achieved by further leveraging temporal locality in ChronosPipe?
Figure 9. End-to-End Evaluation on ChronosPipe. (a) shows the memory saving of different schedules on a 48-layer model at (PP, TP)=(8,8), global batch=128, micro batch=2, seq len=4K. *: R=66.7% in interleave 1F1B, R=50% in 1F1B and Chronos-Recomp is used in Chronos-Pipe. (b) illustrates the maximum trainable model size of different schedules at (PP, TP)=(8,8), global batch=128, micro batch=2, seq len=4K. *: compared with 1F1B.
6.1. Methodology
Setup. Our experiments were carried out on a cluster comprising up to 64 GPUs, organized across 8 nodes. Nodes are interconnected with four 200 Gbps NICs. Each node is equipped with 8 GPUs and 2 CPUs(Intel Xeon Platinum 8480+) and 2TB DRAM, 15.36TB NVME. Each GPU features 32GB of HBM and connects to the CPU through PCIe5 x8(32GB/s), with high-bandwidth interconnects linking GPUs within a node.
Workloads. For these experiments, we utilized a model architecture similar to LLAMA2-70B (LLAMA2, ), consisting of 80 transformer layers and incorporating the GQA mechanism. We varied the number of transformer layers in the model to adjust the size for different experiments.
We further analyze the effectiveness of ChronosPipe on other models through theoretical analysis. Qwen2.5-32B (Qwen, ), PaLM-62B (PaLM, ), and OPT-66B (OPT, ) are selected as they have distinct vocabulary sizes, attention mechanism configurations, and activation functions, being typical enough for current dense models.
Baselines. We compare ChronosPipe against 1F1B and interleave-1F1B to highlight the benefits of introducing temporal locality into pipeline parallelism. All scheduling strategies were implemented within a framework similar to Megatron-LM to ensure a fair comparison. In ChronosPipe’s implementation, only essential asynchronous communications were employed, while all other P2P communications followed the synchronous approach used in the baselines. When recomputation is included, we denote it using ”scheduling strategy + recomputation ratio.” For instance, $1F1B+R=50\%$ represents the 1F1B with a 50% recomputation ratio.
Additionally, through theoretical analysis, we further compare our approach with the recomputation strategies of Megatron-Kwai (MegatronKwai, ) and AdaPipe (AdaPipe, ). Conditions are aligned with AdaPipe’s publicly - available data for a fair comparison.
Unless otherwise specified, the input of ColumnLinear is not evenly divided along the TP dimension, consistent with Megatron’s settings. Meanwhile, Operator-level recomputation (RMSNorm, activation function) and FlashAttention (Flashattention, ) are enabled by default. In theoretical calculations, $T_{c}=0.104T_{unit}$ is set.
Theoretical analysis. Based on the following two reasons, we assert that theoretical analysis demonstrates sufficient credibility in computational and storage evaluation. First, the performance of PP can be effectively modeled. By precisely quantifying the bubble ratio and recomputation proportion, accurate performance estimation becomes achievable—an approach validated as feasible in  (PipeFisher, ; Chimera, ). Regarding storage, Activation and Model State dominate storage consumption, enabling theoretical analysis to effectively identify critical storage constraints.
6.2. End-to-End Evaluation
Fig.  9(a) illustrates that ChronosPipe significantly reduces HBM memory requirements while maintaining comparable throughput. Under the PP8_TP8 configuration, ChronosPipe’s optimized scheduling enables training a 48-layer (i.e., 42B) model, while Interleave-1F1B and 1F1B run into out-of-memory (OOM) issues. Chronos-Recomp further demonstrates improved recomputation efficiency, reducing activation memory from 17.23 GB to 6.63 GB. In contrast, under the same recomputation budget (R=50%), the 1F1B method still occupies 10.6 GB for activations. With Chronos-Offload, ChronosPipe ALL reduces the total memory footprint for activations and model states to 64.6% of what is used in the $1F1B+R=50\%$ setup while achieving 97.58% of its throughput.
Fig.  9(b) explicitly evaluates the effect of ChronosPipe in terms of the maximum size of the trainable model. With limited 32GB HBM, 1F1B can train only a 40-layer model under the PP8_TP8 configuration, while training a 64-layer model requires a 50% recomputation ratio. In contrast, ChronosPipe can train a 48-layer model, and with the Chronos-Recomp strategy, this can be extended to 80 layers. Further leveraging Chronos-Offload enables training even a 96-layer model, all while keeping the max reserved memory at roughly the same level. As a result, ChronosPipe increases the scale of trainable models to 2.4 times the 1F1B, and 1.5 times that of the 1F1B+R=50% configuration. Note that this only shows the advantage of introducing temporal locality into PP. ChronosPipe can further integrate other compatible optimization strategies to optimize HBM usage, which is not reflected in this experiment.
Figure 10. Estimated memory saving for different recomputation-based solutions on GPT3-175B at (PP,TP)=(8,8), global batch=32, micro batch=1, seq len=16K. sequence parallelism (SP) (SeqParallel, ) is used in all and operator-level recomputation(RMSNorm, activation function) is not used in ChronosPipe. *:1F1B with only Operator-level recomputation solution is used. activation-offloading is not under consideration due to it suitable for other solutions.
Fig.  10 compares ChronosPipe with other recomputation solutions. Megatron-Kwai’s Operator-aware recomputation reduces storage by 1.27x with negligible performance loss. Applying Operator-aware recomputation to remaining activations shows diminishing returns, but co-designing PP with recomputation can enhance its benefits. AdaPipe adopts Stage-aware task and recomputation budget allocation across pipeline stages, achieving 1.76x storage savings and 1.26x performance improvement over the 1F1B+R = 100% baseline. Without Operator-aware recomputation on activations of deeper layers, Chronos-Pipe+Chronos-Recomp attains 1.72x storage savings and 1.17x performance gain relative to the same baseline. ChronosPipe ALL further achieves 2.22x storage savings, highlighting the efficiency of Locality-aware design. While Chronos-Pipe+Chronos-Recomp exhibits moderately reduced throughput relative to AdaPipe, it provides a general and simple solution, and enhance adaptability to a variety of changes including model scale, structures, and parallelism configurations. In contrast, AdaPipe necessitates sophisticated search mechanisms and programmatic overhead to implement stage-aware task and recomputation budget allocation, which is a hard integer partitioning problem. Maintaining workload equilibrium across varying constraints mentioned while preserving throughput remains questionable. Moreover, these approaches embody distinct design philosophies, revealing synergistic potential for future integration.
Figure 11. memory and throughput of different schedules under various sequence lengths on a 32-layer model at (DP, PP, TP)=(2,4,8), global batch=128, micro batch=2. *: R=50%, Chronos-Recomp is used in ChronosPipe. **: Chronos-Offload is further adopted.
We further study the behavior of different scheduling strategies across various sequence lengths, shown in Fig. 11. Under the DP2 (ZeRO-1)_PP4_TP8 configuration, Interleave-1F1B exhibits the worst memory efficiency, running out of memory (OOM) at a sequence length of 8k, even with R=50%. Theoretical analysis suggests that smaller PP reduces ChronosPipe’s memory efficiency, resulting in Chronos-Pipe and Chronos-Recomp saving only 12.5% and 25% of activation respectively compared to 1F1B and its recomputation variants in this setup. However, as sequence lengths increase, ChronosPipe’s memory savings become increasingly pronounced, demonstrating superior memory efficiency for longer sequences. Additionally, Chronos-Offload further enhances ChronosPipe’s memory efficiency by reducing model state memory, allowing ChronosPipe to excel as both sequence length and model size scale up. In terms of throughput, while Chronos-Pipe experiences a relative throughput drop of 6%–9% compared to 1F1B, both Chronos-Recomp, and ChronosPipe ALL maintain throughput comparable to 1F1B+R=50% across all sequence lengths, even when throughput across sequence lengths are slightly influenced by Operator Library.
Figure 12. estimated memory for different models at (PP,TP)=(8,8), global batch=128, micro batch=2. Ideal Computation Fraction is defined with following equation:(1- bubble overhead - recomputation overhead).
As shown in Fig.  12, we further evaluated the performance of ChronosPipe on dense models under multiple different parameter configurations. Chronos-Pipe+Chronos-Recomp achieves a storage reduction of 1.21x-1.26x while maintaining a throughput similar to that of 1F1B+R=50%. This enables the training of PaLM-62B and OPT-66B models within the 32GB HBM constraint. Moreover, ChronosPipe ALL achieves a storage reduction of 1.56x-1.58x compared to 1F1B+R=50%. Compared to 1F1B+R=100%, ChronosPipe not only boosts the throughput by approximately 1.15x but also reduces the storage by 1.04x-1.10x.
6.3. Further Analysis of ChronosPipe
Figure 13. analysis on P2P overhead on 24-layer model at (PP,TP)=(4,8), micro batch=2, seq len=4K. Ideal Computation Fraction is defined with following equation:(1- bubble overhead - recomputation overhead).
As illustrated in Fig.  13, we evaluated the impact of P2P communication on ChronosPipe under a PP4_TP8 configuration with a 24-layer model. During execution, Chronos-Pipe showed approximately a 6% lower ideal computation fraction compared to 1F1B. The theoretical analysis attributes 5% of this reduction to the overhead introduced by P2P communication, indicating that the primary gap stems from the additional round of P2P communication. This significant P2P overhead can be attributed to low P2P bandwidth and frequent P2P communication(one P2P communication happens after every three layers of computation in this setup), leading to $T_{c}\approx 0.104T_{unit}$. Despite this, Chronos-Recomp achieves an ideal computation fraction comparable to 1F1B+R=50% ($\leq 3\%$), suggesting that its overall P2P cost is similar to that of 1F1B. Specifically, Chronos-Recomp overlaps roughly half of the forward pass P2P communication with computation. However, due to the limited number of layers in the workload, gradient P2P communication during the backward pass (in chunk 2) does not overlap effectively with recomputation, resulting in minor degradation. In practice, ChronosPipe’s superior memory efficiency allows it to train larger workloads, naturally reducing the relative impact of P2P communication. As the workload scales, Chronos-Recomp achieves better overlap for gradient P2P communication during the backward pass, further mitigating its impact on performance.
Figure 14. evaluate the scalability of Chronos-Offload on a 16-layer model at global batch=128, micro batch=2.
We further verify the scalability of Chronos-Offload, as illustrated in Fig.  14. When training a 16-layer model with a sequence length of 4K on PP4_TP8 setup, Chronos-Pipe achieves only partial overlap (45.45%) between the bubble created during the cooldown phase and the time required for weight gradient offloading and CPU-based optimizer updates. This limited overlap results in a slight increase in overall training time. However, when either the pipeline stage or the sequence length is doubled, the overlap improves significantly to 94.55% and 100%, respectively. Furthermore, Chronos-Offload enables only half of the optimizers to be updated on GPU, leading to a slight decrease in overall training time. These results highlight the scalability of Chronos-Offload, demonstrating its increasing effectiveness as the PP stage and sequence length grows, making it a viable solution for handling larger workloads.
6.4. Further Design Space Exploration
Figure 15. in-depth analysis of Chronos-Recomp. Additional memory required for recomputation is ignored for better illustration.
Fig. 15 illustrates our in-depth analysis of Chronos-Recomp. We theoretically evaluated Chronos-Recomp performance under PP4_TP8 configuration with chunk sizes of 2, 3, and 4. While small PP is less favorable for larger chunks in Chronos-Pipe, a clear trend is observed: Chronos-Recomp prioritizes recomputation tasks based on their temporal locality. When the recomputation budget is limited, focusing recomputation on shallower layers provides higher gain. For example, with a chunk size of 4, recomputing only 25% of the layers can reduce activation memory usage by up to 43.75%. However, as the recomputation ratio increases, Chronos-Recomp incurs a higher cost for recomputation in layers where it is less efficient. Nevertheless, Chronos-Recomp consistently outperforms standard recomputation strategies, as shown in Fig. 15.
Figure 16. in-depth analysis of Chronos-Offload.
When partitioning the workload into more chunks, Chronos-Offload has the potential to offload more model states. Our theoretical analysis, illustrated in Fig.  16, explores this behavior while maintaining throughput. The results indicate that Chronos-Offload experiences diminishing returns as the number of chunks increases; in some cases, memory efficiency can even decrease. This effect arises because adding more chunks only marginally increases the number of usable bubbles in the cooldown phase, particularly under smaller PP configurations. For example, when the PP stage is set to 4, using chunk=3 only provides 50% more bubbles compared to chunk =2, while chunk =4 produces the same number of bubbles as chunk =3. Consequently, the advantages of chunk =3 are mainly evident with long sequence length, whereas chunk =4 underperforms because the number of chunks matches the PP stage, ultimately reducing overall memory efficiency.
7. RELATED WORK
Hybrid Parallelism. To address the challenges of scaling model size and sequence length, modern distributed training systems leverage GPU clusters to expand overall memory capacity while deploying various schedules. These strategies include data parallelism(DP), pipeline parallelism(PP), tensor parallelism(TP), context parallelism(CP), and so on. DP (ZeRO, ; FSDP, ), PP (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ), and TP (AccPar, ; 2D-TP, ; 2p5D-TP, ; 3D-TP, ) enable the distribution of model states across machines. CP (RingAttention, ; Ulysses, ) and sequence parallelism(SP) (SeqParallel, ) specialize in distributing activations across their respective dimensions. Hybrid parallelism combines these approaches, enabling efficient scaling to models with trillions of parameters. PP is an essential component of hybrid parallelism among these strategies, particularly in cross-node deployments, due to its low communication requirement. However, PP suffers from pipeline bubbles and high activation memory overhead. Moreover, it is often less compatible with advanced DP like ZeRO-2/3.
Pipeline Parallelism. Most pipeline parallelism research focuses on reducing pipeline bubbles by finer-grained task partitioning  (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ) or multi-dimensional task scheduling  (Chimera, ; Mixpipe, ). However, they neglect the significant overhead of activation memory.  (CtrlMemPP, ) reduces activation lifespans through scheduling but risks introducing more bubbles and prolonging gradient lifespan. It’s V-shaped scheduling also complicates compatibility with ZeRO-2/3 and Chronos-Offload, making storage optimization for other data more challenging.  (AdaPipe, ) addresses activation memory imbalance with asymmetric partitioning but increases programming complexity.  (BreadthFirstPP, ; ZeroPP, ), achieve ZeRO-2/3-compatibile PP at the cost of higher activation memory.  (LLAMA3, ) takes advantage of  (BreadthFirstPP, ) and 1F1B for improved ZeRO-2 compatibility but only maintains, rather than reduces, activation memory usage.
Recomputation and Offloading. In addition to optimizing multi-GPU systems’ HBM utilization through hybrid parallelism, memory requirements can also be reduced by trading memory for computation or communication time. For recomputation, some studies  (SeqParallel, ; MegatronKwai, ; SuperNeuron, ; CapuChin, ) focus on Selective recomputation of different operators but fail to address the high memory usage of the projection operator.  (DTR, ) introduced temporal locality-aware recomputation principle and demonstrated its efficacy in DP, but coarse-grained PP and risk of inducing steady-phase bubbles limit the application of this principle in PP. Some approaches leverage the characteristics of Hybrid parallelism.  (AdaPipe, ) adapts recomputation to mitigate activation memory imbalance in PP.  (OverlapRecomp, ) overlaps recomputation with collective communication in the TP dimension to reduce overhead. However, these approaches increase programming complexity. For offloading,  (Bpipe, ) addresses memory imbalance in PP by offloading Activation across HBM. Most studies focus on fully utilizing storage systems. For example,  (ZeRO-Offload, ; PatrickStar, ; Mobius, ) explore offloading activation and model states to DRAM, while  (ZeRO-Infinity, ; Angel-ptm, ; Smart-Infinity, ; Fuyou, ) extend this to NVMe. Due to limited bandwidth in offloading, these efforts primarily focus on identifying opportunities to overlap offloading with computation and realize near-data processing.
8. CONCLUSION
This work introduces ChronosPipe to address the memory capacity limitations in LLM training. The core idea behind ChronosPipe is to treat HBM as a fast but limited ”cache,” optimizing and leveraging temporal locality in LLM pretraining for efficient HBM usage. Chronos-Pipe enhances activation temporal locality in PP scheduling and, together with Chronos-Recomp and Chronos-Offload, discards data with poor temporal locality from HBM. Experiments show that ChronosPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation. At the same time, ChronosPipe has good enough compatibility to make it suitable for large-scale hybrid parallelism pre-training.
Appendix A Formalization on Delay Rounds at Chronos-Recomp
In Chronos-Pipe scheduling, an additional interval time $T_{fwd\_interval}$ exists between forward chunks, derived from steps 1, 2, and 3 in Fig. 5(a), resulting in $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$. In Chronos-Recomp’s shallow-layer full recomputation, the execution times for steps 1, 2, and 3 respectively change $T_{unit},\lceil\frac{(p-3)}{6}\rceil T_{unit}$, and $\lceil\frac{(p-1)}{2}\rceil T_{unit}$. Therefore, $\Delta T_{fwd\_interval}$ becomes $(\lceil\frac{(p-1)}{2}\rceil-\lceil\frac{(p-3)}{6}\rceil-1)T_{unit}(P\geq 3)$. Based on this, We can formalize the number of rounds k required to delay the launch of chunk 2 as an optimization problem for $P\geq 3$:
(8) $Min$ $\quad k$
$s.t.\quad T_{fwd\_interval}-$ $\Delta T_{fwd\_interval}+7kT_{unit}\geq 0$
$k$ $\in N$
$p$ $\geq 3$
Therefore, when $8\leq p\leq 40$, we have $k=1$, meaning that chunk 2 needs to be delayed by one round.

### v2

@@
 Abstract.
-Larger model sizes and longer sequence lengths have empowered the Large Language Model (LLM) to achieve outstanding performance across various domains.
-However, this progress brings significant storage capacity challenges for LLM pretraining.
-High Bandwidth Memory (HBM) is expensive and requires more advanced packaging technologies for capacity expansion, creating an urgent need for memory-efficient scheduling strategies.
-Yet, prior pipeline parallelism schedules have primarily focused on reducing bubble overhead, often neglecting memory efficiency and lacking compatibility with other memory-efficient strategies.
-Consequently, these methods struggle to meet the storage demands of storage capacity for next-generation LLM.
-This work presents ChronosPipe, a Chronos-aware pipeline parallelism for memory-efficient LLM pretraining.
-The core insight of ChronosPipe is to treat HBM as a fast but small ’cache,’ optimizing and exploiting temporal locality within LLM pretraining to enhance HBM utilization.
-ChronosPipe introduces a pipeline scheduling strategy, Chronos-Pipe, to reduce the extrinsic overhead that disrupts the temporal locality of activations.
-Additionally, it leverages Chronos-Recomp and Chronos-Offload to efficiently harness the intrinsic temporal locality of activations and weights in Deep Neural Networks.
-Experiment results show that ChronosPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
+Scaling LLMs through increased parameters and sequence lengths imposes significant storage pressure on LLM training.
+While expensive HBM requires advanced packaging for capacity scaling, existing pipeline parallelism strategies prioritize computational bubble reduction over memory efficiency.
+Consequently, existing methods fail to address the increasing storage requirements for ever-growing LLMs.
+This work presents TPipe, a temporal-locality-aware pipeline parallelism for memory-efficient LLM training.
+The core insight of TPipe is to treat HBM as a fast but small ’cache,’ optimizing and exploiting temporal locality within LLM.
+TPipe solution comprises three components: T-Pipe, T-Recomp, and T-Offload.
+T-Pipe, a pipeline scheduling strategy, is used to reduce the extrinsic overhead that disrupts the temporal locality of activations.
+T-Recomp and T-Offload are utilized to efficiently harness the intrinsic temporal locality of the activations and weights in DNN.
+Experiment results show that TPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
 Pipeline Parallelism, Recomputation, Offload, Memory-Efficient Schedule, Temporal Locality
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
 1. INTRODUCTION
@@
 For instance, HBM accounts for 50-60% of the total costs in NVIDIA’s H100 SXM5 module (H100, ).
-Moreover, next-generation HBM poses considerable problems.
-Increasing the count of stacking DRAM die within the restricted packaging height (720 um) necessitates hybrid bonding (HybridBonding, ), whose thermal management and yield optimization (HybridBonding_Problem, ) are extremely challenging.
+Moreover, increasing the count of stacking DRAM die within the restricted packaging height (720 um) necessitates hybrid bonding (HybridBonding, ), whose yield optimization (HybridBonding_Problem, ) is extremely challenging.
 Therefore, LLM training necessitates memory-efficient scheduling strategies while existing Pipeline Parallelism (PP) predominantly emphasizes performance optimization and overlooks memory constraints.
@@
 However, unlike the offloading of optimizer states, which only happens once per mini-batch, activation offloading is more frequently conducted across different micro-batches, thus raising a higher demand for offloading bandwidth.
-Figure 1. (a) limited HBM capacity is difficult to meet the demands of next-generation LLM. (b) existing pipeline parallelism ignores temporal locality
-Moreover, current LLM training predominantly employs hybrid parallelism, yet existing PP strategies lack optimal compatibility with other memory-efficient scheduling approaches.
-The ZeRO series (ZeRO, ) represents state-of-the-art data parallelism (DP) techniques, distributing model states across DP machines.
-In hybrid DP and PP configurations, distributing optimizer states (ZeRO-1) necessitates inter-DP rank communication only at mini-batch granularity.
-However, further distributing weight gradients (ZeRO-2) and weights (ZeRO-3) requires more frequent communication at micro-batch granularity, imposing higher bandwidth demands between DP ranks.
-Consequently, hybrid DP-PP implementations (DeepSpeed, ) typically utilize only ZeRO-1, leaving substantial potential for reducing weight and gradient storage.
-When designing PP strategies that are more compatible with ZeRO-2/3, such as Breadth First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), these approaches reduce DP bandwidth requirements at the cost of a significant increase in activation memory, leaving no perfect solution.
+Figure 1. (a) limited HBM capacity is difficult to meet the demands of next-generation LLM. (b) existing pipeline parallelism (such as 1F1B) ignores temporal locality.
 Existing PP strategies lack memory efficiency due to their failure to address an intrinsic property of Deep Neural Network (DNN): temporal locality.
@@
 Unlike previous work, we notice the issues mentioned above and introduce the concept of temporal locality to pipeline parallelism for the first time.
-We propose ChronosPipe, which enables the training of models 2.4 times larger than 1F1B (DAPPLE, ) on the same hardware with comparable throughput while maintaining compatibility with various memory-efficient scheduling strategies.
-The ChronosPipe solution comprises three components: Chronos-Pipe, Chronos-Recomp, and Chronos-Offload, which are described below.
-Chronos-Pipe minimizes peak activation storage by reducing micro-batch execution time, thus expediting the backward pass that consumes activations.
+We propose TPipe, which enables the training of models 2.4 times larger than 1F1B (DAPPLE, ) on the same hardware with comparable throughput while maintaining compatibility with various memory-efficient scheduling strategies.
+The TPipe solution comprises three components: T-Pipe, T-Recomp, and T-Offload.
+T-Pipe minimizes peak activation storage by reducing micro-batch execution time, thus expediting the backward pass that consumes activations.
 This, in effect, optimizes activation temporal locality.
-Chronos-Recomp exploits the poorer temporal locality of shallower layer activations and discards them from HBM by selectively recomputing to achieve higher efficiency.
-Chronos-Offload takes advantage of poorer temporal locality of deeper layer weight and discards them from HBM by offloading the process of optimizer update to CPU.
-Since this process overlaps with PP’s warmup and cooldown phase, it lowers demand on offload bandwidth and the processing speed of CPU.
+T-Recomp exploits the poorer temporal locality of shallower layer activations and discards them from HBM by selectively recomputing to achieve higher efficiency.
+T-Offload takes advantage of poorer temporal locality of deeper layer weight by offloading optimizer update to CPU during bubbles created in T-Pipe.
+Since this process overlaps with warmup/cooldown phase, it lowers demand on offload bandwidth and CPUs.
+These strategies co-design based on temporal locality variations in pipeline parallelism, which remain underexplored in existing methods.
 Our main contributions are as follows.
-  • We found that temporal locality in DNN is the primary cause of imbalanced activation storage in PP. This imbalance limits the size of trained models on current devices.
-  • For activation memory savings, we introduce temporal locality into the PP schedule (Chronos-Pipe) and recomputation (Chronos-Recomp).
-Chronos-Pipe eliminates unnecessary intervals, optimizing the temporal locality of activation (Section 4.1).
-Chronos-Recomp discards activation with poor temporal locality from HBM by selectively recomputing shallow layers to achieve higher efficiency (Section 4.2).
-  • For Model State storage capacity, we show that a ZeRO-2-compatible PP could built based on Chronos-Pipe (Section 4.3).
-Additionally, we designed the offload strategy based on temporal locality (Chronos-Offload), which can discard model states with poor temporal locality from HBM by offloading optimizer updates of deeper layers to the CPU (Section 5.1).
-  • End-to-end evaluation on ChronosPipe is carried out on a cluster comprising up to 64 accelerators.
-Experiments show that Chronos-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
+  • We found that temporal locality in DNN is the primary cause of imbalanced activation storage in PP, constraining trainable model sizes on current devices.
+  • For activation memory savings, we introduce temporal locality into the PP schedule (T-Pipe) and recomputation (T-Recomp).
+T-Pipe eliminates unnecessary intervals, optimizing the temporal locality of activation (Section 4.1).
+T-Recomp discards activation with poor temporal locality from HBM by selectively recomputing to achieve higher efficiency (Section 4.2).
+  • For model state storage, we designed the offload strategy based on temporal locality (T-Offload), which can discard model states with poor temporal locality from HBM by offloading optimizer updates of deeper layers to CPU during bubbles created in T-Pipe (Section 5).
+  • End-to-end evaluation on TPipe is carried out on a cluster comprising up to 64 GPUs.
+Experiments show that T-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
 2. BACKGROUND AND MOTIVATION
-In this section, we will introduce pipeline parallelism (PP) (Section 2.1), typical recomputation strategies (Section 2.2), and offloading strategies (Section 2.3).
-We will then discuss the motivation for leveraging temporal locality (Section 2.4).
+In this section, we will introduce pipeline parallelism (PP) (Section 2.1), typical recomputation strategies and offloading strategies (Section 2.2).
+We will then discuss the motivation for leveraging temporal locality (Section 2.3) and review how prior work has addressed memory imbalance challenges in PP (Section 2.4).
 To facilitate clearer explanations throughout the paper, we will use the following symbols.
 $\bullet$ $m_{a}$:memory consumption for activation in whole Neural Network(exclude input embedding layer and language modeling head)
-$\bullet$ $P$:# of pipeline stages
-$\bullet$ $m$:# of micro-batches executed within a training iteration
-$\bullet$ $T_{fwd}$: execution time of forward pass in one micro-batch
-$\bullet$ $T_{bwd}$: execution time of backward pass in one micro-batch
+$\bullet$ PP:# of pipeline stages
+$\bullet$ mm:# of micro-batches within a training iteration
+$\bullet$ $T_{fwd}/T_{bwd}$: execution time of forward/backward pass in one micro-batch
 2.1. Pipeline parallelism
-Figure 2. (a) Interleave-1F1B schedule. (b) ZeRO-compatible pipeline schedule
+Figure 2. (a) Interleave-1F1B (b) and its memory usage
 Pipeline Parallelism (PP) divides the model into multi-layer blocks mapped across different PP stages.
 These stages work together to train all samples within a mini-batch, enabling mini-batch Stochastic Gradient Descent.
-During the forward pass, samples in each mini-batch begin computation at the first stage, with subsequent blocks executing their computations in a pipelined manner until the final forward computation completes at the last stage.
-In the backward pass, gradients for each mini-batch flow in the reverse direction.
-As a result, PP only requires a single tensor to be sent and received per block, significantly reducing communication bandwidth demands.
-This approach is often deployed across nodes and is a key component in hybrid parallelism.
-Pipeline parallelism (PP) at the mini-batch level suffers from significant pipeline bubbles, prompting the common practice of splitting mini-batches into smaller micro-batches.
-By dividing the workload, the first stage, after processing the initial micro-batch, can immediately start on subsequent micro-batches rather than idling, thus reducing pipeline bubbles (GPipe, ).
-When the number of micro-batches increases, the classic 1F1B (one-forward-one-backward) scheduling strategy (DAPPLE, ) is often used, as shown in Fig.
+Forward computations in each mini-batch propagate sequentially through stages until completion, while gradients flow inversely during backward passes.
+As a result, only one layer of intermediate activation/gradient needs to be sent and received per block, significantly reducing communication bandwidth demands.
+Mini-batch-level PP suffers from significant pipeline bubbles, commonly addressed by splitting mini-batches into micro-batches.
+Dividing workloads lets the first stage proceed from initial to subsequent micro-batches without idling, thus reducing pipeline bubbles (GPipe, ).
+When the number of micro-batches increases, the classic 1F1B strategy (DAPPLE, ) is often used, as shown in Fig.
 1(b).
-During the steady phase, a forward computation block is initiated only after a backward block finishes, avoiding the simultaneous launch of multiple forward passes in the warm-up phase, which would otherwise increase peak memory usage.
+During the steady phase, each forward computation block follows a backward block completion, preventing concurrent forward passes during warm-up that elevate peak memory usage.
 However, this approach creates an imbalance in activation memory usage across PP stages.
-In a single micro-batch, the first stage starts forward computation early but is the last to complete backward passes and release memory.
+In a single micro-batch, the first stage initiates forward computation first but finalizes backward passes and release memory last.
 As a result, the first stage reaches peak activation memory of $m_{a}$, while the last stage’s activation memory usage is only $\frac{m_{a}}{p}$.
-Additionally, Fig.
+Fig.
 1(b) illustrates that tasks with dependencies between adjacent PP stages execute sequentially, leading to non-overlapping point-to-point (P2P) communication.
-Some scheduling methods address this by inserting independent task blocks from other micro-batches between dependent blocks, enabling overlapping P2P communication.
-This delay caused by dependency constraints is called the ”interval” delay in this paper.
+Some methods mitigate this by inserting independent tasks from other micro-batches between dependent blocks to enable communication overlap.
+This delay caused by dependency is called the ”interval” in this paper.
 Blocks can be broken down into smaller tasks to reduce pipeline bubbles further, though this may increase memory requirements.
-The Interleaved-1F1B approach (Interleave-1F1B, ) divides blocks into more chunks and schedules multiple chunks launches during the warmup phase, as shown in Fig.
+Interleaved-1F1B  (Interleave-1F1B, ) divides blocks into more chunks and schedules multiple chunks launches during the warmup phase, as shown in Fig.
 2(a).
-When the workload is divided into v chunks (assigning layer 1, $p+1$, …, $(v-1)p+1$ to stage 0; layer 2, $p+2$, …, $(v-1)p+2$ to stage 1, and so forth), the bubble size is reduced to $\frac{1}{v}$ of that in the traditional 1F1B, but peak activation memory increases to $m_{a}(1+\frac{p-1}{pv})$.
-The Zero Bubble approach (ZeroBubble, ) goes a step further by splitting the backward pass (BP) into two phases: activation gradient computation (BPA) and weight gradient computation (BPW).
-This arrangement theoretically enables a bubble-free pipeline.
-However, the activation gradients generated by BPA are required to be retained until BPW is completed, which raises memory demands.
-Additionally, splitting the BP disrupts the original reuse of activation gradients for all-gather operations between BPA and BPW, and reduces the overlap between computation and communication.
-Designing a ZeRO-compatible pipeline schedule is challenging due to the trade-off between DP communication bandwidth and activation memory storage.
-Communication between DP ranks at the micro-batch level is required for a schedule to be compatible with ZeRO-2 or ZeRO-3.
-To overlap this communication, existing methods, such as Breadth-First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), use grouped execution of micro-batches, as shown in Fig.
-2(b).
-This approach processes multiple micro-batches simultaneously, allowing communications of weights and weight gradients can be overlapped with computations, but it also causes a substantial increase in activation memory usage.
-Since weights and weight gradients are stored in 16-bit and 32-bit precision, respectively, achieving ZeRO-3 PP requires at least a 1.5x increase in DP communication bandwidth or an increase in the number of micro-batches per group compared to ZeRO-2 PP.
-Consequently, some implementations choose to use only ZeRO-2 PP (LLAMA3, ).
-2.2. Recomputation
+When the workload is divided into v chunks (assigning layer 1, $p+1$, …, (v1)p+1(v-1)p+1 to stage 0; layer 2, $p+2$, …, (v1)p+2(v-1)p+2 to stage 1, and so forth), the bubble size is reduced to $\frac{1}{v}$ of that in the traditional 1F1B, but peak activation memory increases to $m_{a}(1+\frac{p-1}{pv})$ (Fig.
+2(b)).
+Zero Bubble (ZeroBubble, ) splits backward propagation (BP) into activation gradient (BPA) and weight gradient (BPW), enabling a nearly ’bubble-free’ pipeline.
+BPA-generated gradients must persist until BPW completion, which raises memory demands.
+Additionally, splitting the BP disrupts reuse of activation gradients for all-gather operations between BPA and BPW, and reduces the overlap between computation and communication.
+2.2. Recomputation and Offloading Strategy
 Recomputation allows only a subset of activations generated in the forward pass (called checkpoints) to be retained, with all other activations regenerated during the backward pass, trading off memory usage for additional computation.
-Previous work (ChainTrainingMem, ) has explored optimal checkpoint placement, given a fixed number of checkpoints, to minimize activation memory, achieving memory savings at sub-linear cost for chain structure of DNN.
+Previous work (ChainTrainingMem, ) introduced recomputation into DNNs, proving that optimal scheduling enables memory savings at sublinear cost for chain-structured DNN.
 Additionally, the benefits of recomputation vary across different operators.
@@
 Moreover,  (DTR, ) further introduced temporal locality-aware recomputation principle and demonstrated its efficacy in data parallelism.
-Although pipeline stages inherently exhibit temporal locality variations in PP, coarse-grained PP fails to leverage intra-block temporal locality variance for recomputation optimization within the simplistic structure of LLMs.
-Existing methods thus fail in single-chunk PP.
-Moreover, Direct application of this principle to multi-chunk PP (e.g., interleaved 1F1B scheduling) risks inducing steady-phase bubbles, resulting in significant throughput degradation.Therefore,  (DTR, ) fail to solve imbalanced memory in PP.
-2.3. Offloading Strategy
+However, temporal locality-aware recomputation in pipeline parallelism remains underexplored, necessitating systematic solutions for inter-stage dependency management and programming complexity mitigation.
 Offloading techniques allow tensors to be moved from memory-constrained GPUs to other GPUs or CPUs, trading memory usage for data movement.
 In BPipe (Bpipe, ), GPU offloading is used to address activation memory imbalances in pipeline parallelism.
-However, this approach requires high-bandwidth intra-node communication links like NVLink rather than slower inter-node links like Ethernet or InfiniBand.
+However, this approach requires high-bandwidth intra-node communication links like NVLink.
 CPU offloading utilizes PCIe to offload activation or model states.
 Theoretically, CPU offloading for activation memory reduction is well-suited for long-sequence training (MegatronKwai, ).
-However, limited PCIe bandwidth may leave some activation in HBM in practice.
-For model states, previous work, such as ZeRO-Offload (ZeRO-Offload, ), offloads the process of optimizer-step to the CPU, storing both optimizer states and master weights in CPU-side DRAM.
+However, completely offloading all the activations can lead to significant performance degradation due to limited PCIe bandwidth.
+For model states, previous work, such as ZeRO-Offload (ZeRO-Offload, ), offloads optimizer-step to CPUs.
 To minimize GPU idle time caused by latency in the CPU’s optimizer step, this approach imposes significant requirements on both offload bandwidth and CPU computational capacity.
 Overall, Recent studies still have high bandwidth requirements for the Offloading Strategy.
-2.4. Motivation: Temporal-Locality Matters
+2.3. Motivation 1: Temporal-Locality Matters
 Figure 3. forward and backward pass of deep neural network
@@
 Higher Peak Activation Storage Requirements: The PP stage responsible for shallow-layer activations releases them last, which increases the activation storage requirements.
-Additionally, intervals can exist between blocks with dependencies, and these intervals can accumulate across multiple blocks, further worsening the temporal locality for shallow layers.
-This is the reason that the interleaved 1F1B schedule results in a higher peak storage demand.
+Additionally, intervals can exist between blocks with dependencies, and these intervals can accumulate across multiple blocks, further worsening the temporal locality.
 Skewed Distribution of Peak Activation Storage: As shown in Fig.
-2(a), the peak activation storage in the interleaved 1F1B schedule tends to occur at Stage 0.
-When the chunk size is set to 2, the ratio of peak activation storage between shallow and deep layers is $(2p-1):p$, indicating an obvious bias.
-Sufficient Time for Process of optimizer-step: Beyond the natural advantage of chain-structured networks providing additional time for optimizer update on deep-layer weights, increasing the PP will provide more sufficient time for optimizer-step.
-Large PP does not impact the execution latency of each micro-batch while the amount of model state that needs updating per stage is reduced, thereby lowering the communication and processing demands for weight updates.
-3. OVERVIEW OF ChronosPipe
+2, the peak activation storage in the interleaved 1F1B tends to occur at Stage 0.
+When the chunk size is set to 2, the ratio of peak activation storage between shallow and deep layers is (2p1):p(2p-1):p, indicating an obvious bias.
+Sufficient Time for optimizer-step: Beyond the natural advantage of chain-structured networks providing additional time for optimizer update on deep-layer weights, increasing the PP will provide more sufficient time for optimizer-step.
+Larger pipeline parallelism does not impact execution latency of each micro-batch while reducing amount of per-stage model state updates.
+Additionally, warm-up/cool-down phase bubbles can be strategically utilized to reduce communication and processing overhead during weight updates.
+2.4. Motivation 2: Memory-efficient PP Solutions
+To address the memory imbalance problem in pipeline parallelism (PP), prior work primarily follow two directions.
+The first way employs stage-aware strategies like AdaPipe (AdaPipe, ), which dynamically allocates varying computational workloads and recomputation ratios across pipeline stages to achieve memory balance.
+However, this adaptive allocation introduces significant compilation and programming complexity as it must continuously adjust to different model architectures and hardware configurations.
+The second approach adopts locality-balanced solutions exemplified by V-Half/V-Min  (CtrlMemPP, ), which utilize a V-shaped scheduling pattern to assign tasks with the best temporal locality (deepest layer) and worst temporal locality (shallowest layer) within the same pipeline stage, thus statically balancing memory across stages while maintaining uniform computational loads.
+Despite their merits, V-Half/V-Min exhibit critical limitations.
+First, while achieving stage-level memory balance through V-shaped scheduling, these methods still retain tensors with poor temporal locality, failing to reach the upper bound of memory optimization via temporal locality.
+Second, their memory-efficient V-shaped design requires partitioning backpropagation (BP) into separate BPA and BPW phases, preventing overlap between tensor parallelism’s collective communications and computation.
+This inefficiency becomes particularly severe in long-sequence training, where it introduces persistent pipeline bubbles and degrades computational efficiency.
+Unlike previous work, we propose TPipe, a rigorous temporal-locality-aware design to systematically address these limitations, as detailed in the following sections.
 Figure 4.
-Overall scheduling of ChronosPipe.
-The core idea behind ChronosPipe is to treat HBM as a fast but limited ”cache.” Chronos-Pipe enhances activation temporal locality in PP scheduling while Chronos-Recomp and Chronos-Offload discard activation and weight with poor temporal locality from HBM.
-Inspired by the principle of efficient storage utilization in Cache, ChronosPipe applies cache’s core concepts to HBM.
-The central idea of ChronosPipe is to optimize and leverage temporal locality during LLM pretraining, enabling more efficient utilization of the HBM.
-ChronosPipe classifies the factors affecting temporal locality in LLM training into two main types.
-First, temporal locality varies across different layers within the same network; this influence is referred to as intrinsic temporal overhead, capturing the impact on temporal locality inherently introduced by the model’s structure.
+Overview of TPipe.
+The core idea behind TPipe is to treat HBM as a fast but limited ”cache.” T-Pipe enhances activation temporal locality by scheduling while T-Recomp/T-Offload discard activation/weight with poor temporal locality from HBM.
+3. OVERVIEW OF TPIPE
+Inspired by the principle of locality in cache, TPipe applies cache’s core concepts to HBM.
+The central idea of TPipe is to optimize and leverage temporal locality during LLM training, maximizing HBM utilization efficiency.
+TPipe classifies the factors that affect temporal locality in LLM training into two main types.
+First, temporal locality varies across different layers within the same model; this influence is referred to as intrinsic temporal overhead, capturing the impact on temporal locality inherently introduced by model structure.
 Second, the same layer may display different temporal locality depending on the scheduling strategy used; this influence is termed extrinsic temporal overhead, capturing the effects on temporal locality introduced by scheduling differences.
-To address extrinsic temporal overhead, ChronosPipe draws on the first critical insight from Cache: optimize locality in applications.
+To address extrinsic temporal overhead, TPipe draws on the first critical insight from cache: optimize locality in applications.
 As illustrated in Fig.
-4, ChronosPipe introduces a PP scheduling strategy called Chronos-Pipe to advance the backward passes of both shallow and deep layers, minimizing the lifespan of activation.
-Detailed explanations of Chronos-Pipe are provided in Section 4.1.
-This scheduling strategy reflects ChronosPipe’s approach to optimizing temporal locality in LLM training.
+4, TPipe introduces a PP scheduling strategy called T-Pipe to advance backward passes across all layers, minimizing the lifespan of activation.
+Detailed explanations of T-Pipe are provided in Section 4.1.
+This scheduling strategy reflects TPipe’s approach to optimizing temporal locality in LLM training.
 While scheduling can only reduce extrinsic temporal overhead, addressing intrinsic temporal overhead still requires additional strategies.
-To tackle this problem, ChronosPipe draws on a second critical insight from Cache: discard data with poor locality from Cache.
+To tackle this problem, TPipe draws on a second critical insight from cache: discard data with poor locality from cache.
 As illustrated in Fig.
-4, ChronosPipe removes data with poor intrinsic temporal locality from HBM.
-This involves Chronos-Recomp, a recomputation strategy for shallow-layer activations, and Chronos-Offload, an offloading strategy for deep-layer weights.
-Detailed explanations of Chronos-Recomp and Chronos-Offload are provided in Sections 4.2 and 5.1, respectively.
-Through recomputation and offloading, ChronosPipe leverages temporal locality to improve memory efficiency in LLM training.
-4. Chronos-aware SCHEDULE FOR ACTIVATION MEMORY SAVINGS
-4.1. Chronos-Pipe: Chronos-aware Pipeline
-Figure 5. (a) Chronos-Pipe Pipeline Schedule(chunk size=2). (b) Analysis of overhead of P2P communication in Chronos-Pipe.
-Chronos-Pipe Schedule
-The primary goal of Chronos-Pipe is to mitigate the impact of interval accumulation in pipeline scheduling, which worsens activation temporal locality.
-To leverage the temporal locality differences across layers, we further split blocks into multiple chunks and prioritize completing backpropagation as early as possible.
-Unlike interleaved 1F1B, we aim to execute tasks with dependencies in adjacent stages as tightly as possible to minimize the negative effects of interval accumulation across multiple tasks on shallow-layer temporal locality.
+4, TPipe removes data with poor intrinsic temporal locality from HBM.
+This involves T-Recomp, a recomputation strategy for shallow-layer activations, and T-Offload, an offloading strategy at bubble for deep-layer weights.
+Detailed explanations of T-Recomp and T-Offload are provided in Section 4.2 and Section 5, respectively.
+Through recomputation and offloading, TPipe leverages temporal locality to improve memory efficiency in LLM training.
+4. TEMPORAL-LOCALITY-AWARE SCHEDULE FOR ACTIVATION
+4.1. T-Pipe: Temporal-Locality-Aware Pipeline
+Figure 5. Comparison of Interleave-1F1B, 1F1B, and T-Pipe.
+T-Pipe Schedule
+We first analyze memory usage in 1F1B and Interleave-1F1B through intrinsic and extrinsic temporal overhead perspectives.
+As Fig.
+5 illustrates, while different DNN layers exhibit varying temporal locality, 1F1B’s coarse-grained execution fails to leverage intrinsic locality differences among layers assigned to the same stage.
+Although Interleave-1F1B employs chunk-based fine-grained pipelining to reveal intrinsic locality variations within stages, it incurs higher activation memory than 1F1B due to significant extrinsic temporal overhead.
+For example, the extrinsic overhead caused by accumulated intervals during the computation of chunk 2 (micro-batch 4) equals the intrinsic overhead required for DNN computation.
+Therefore, T-Pipe’s core innovation lies in reducing extrinsic temporal overhead while leveraging intrinsic locality variations through fine-grained scheduling.
+Similar to Interleave-1F1B, T-Pipe partitions workloads into chunks to expose intrinsic locality variations within stages (Fig.
+5).
+However, it uniquely prioritizes early backpropagation completion to minimize extrinsic overhead.
+Interleave-1F1B and T-Pipe can be viewed as dual optimization approaches: while both employ chunk-based fine-grained pipelining, the former prioritizes computational performance, whereas the latter emphasizes memory efficiency.
+Guided by extrinsic overhead minimization principles, T-Pipe achieves a peak activation memory of just 75% $m_{a}$, halving Interleave-1F1B’s activation storage requirements.
+This memory advantage comes with a fundamental trade-off: The bubble ratio of T-Pipe doubles that of Interleave-1F1B, similar to 1F1B.
+These quantitative results are rigorously derived from the theoretical analysis presented in Appendix A.
+Driven by intrinsic temporal locality variations, shallow-layer activations account for 50% $m_{a}$, while deep-layer activations contribute 25% $m_{a}$.
+We will implement targeted technique to exploit this fundamental temporal disparity, as detailed in the subsequent section.
+Figure 6.
+(a) T-Pipe + Standard Recomp(R=50%).
+(b) Dependency conflict would occur in T-Pipe+layer-grouped T-Recomp.
+(c) removing dependency conflict in T-Pipe+layer-grouped T-Recomp results in large extrinsic temporal overhead.
+(d) T-Pipe+block-wise T-Recomp.
+(e) Interleave-1F1B+block-wise T-Recomp.
+4.2. T-Recomp: Temporal-Locality-Aware Recomputation
+Why T-Recomp
+T-Pipe successfully minimizes extrinsic temporal overhead, but remains fundamentally constrained by the intrinsic temporal locality differences in the chain structure of DNN.
+To further improve memory efficiency, we can evict data with poor temporal locality from HBM.
+This aligns with fundamental principle in computer architecture: caches with limited capacity should retain only data exhibiting good locality.
+We systematically adapt this well-established caching paradigm to HBM management.
+To address eviction of low-temporal-locality data from HBM, the most intuitive approach involves offloading target activations to storage in next level (e.g., CPU DRAM).
+However, this solution demonstrates critical limitations: its performance becomes strictly bandwidth-bound due to repeated offload/reload cycles across all training batches, each requiring identical data movement operations.
+Given that compute scaling has consistently outpaced memory and communication bandwidth scaling historically, this approach is nonuniversal: In systems exhibiting significant compute-to-bandwidth asymmetries (e.g., wafer-scale integrated system (Dojo, ) (Cerebras, ) or PCIe-bandwidth-constrained GPU clusters), it incurs severe throughput degradation.
+A more viable solution involves directly evicting low-temporal-locality data from the memory hierarchy and recomputing it when required.
+Given that computation scaling consistently outpaces memory and communication bandwidth scaling in hardware evolution, rendering computational resources relatively inexpensive, computation emerges as a hardware-agnostic technique without presupposing specific memory or interconnect bandwidth configurations.
+We therefore employ recomputation to evict activation values with poor temporal locality from HBM.
+This approach strategically prioritizes recomputation of shallower layers, a methodology we designate as T-Recomp.
+Concerns Regarding T-Recomp
+T-Recomp fundamentally differs from conventional recomputation in pipeline parallelism.
+Traditional coarse-grained PP fails to reveal temporal locality variations within single stage, necessitating T-Recomp’s implementation in multi-chunk fine-grained pipelining.
+Under standard recomputation, all chunks perform recomputation with same ratio, resulting in uniform backpropagation times across chunks.
+However, T-Recomp strategically prioritizes recomputation of shallower-layer activations, intentionally creating nonuniform backpropagation times between chunks, a key innovation that aligns recomputation with intrinsic temporal locality patterns.
+Conventional recomputation employs a layer-grouped strategy for memory efficiency, partitioning transformer layers within each chunk into discrete groups while preserving only input activation of each group.
+During recomputation, buffer must be allocated for regenerated activation tensors.
+To minimize buffer capacity, recomputation is triggered just-in-time when activations of specific group are required.
+This design prevents accumulation of recomputed results, which has become prevalent in standard implementations.
+Unfortunately, the layer-grouped recomputation approach is incompatible with T-Recomp.
+This method forces subsequent computations within the same stage to wait not only for the current recomputation but also for accumulated delays from preceding stages.
+While conventional recomputation maintains uniform backpropagation times across chunks, preserving consistent inter-stage dependencies, T-Recomp’s variable recomputation times (as shown in Fig.
+6(b)) disrupt this balance.
+The resulting accumulation of recomputation alters inter-stage dependency patterns, creating dependency conflicts.
+Attempting to resolve these violations through schedule adjustments like advancing conflicting tasks as illustrated in Fig.
+6(c) introduces substantial extrinsic temporal overhead, ultimately undermining motivation for temporal locality optimization.
+Addressing the Concerns
+The key solution lies in preventing accumulated recomputation from affecting dependency relationships, which we achieve through block-wise T-Recomp.
+As illustrated in Fig.
+6(d), we decouple recomputation from the dependent backpropagation process, thereby eliminating the negative impact of recomputation accumulation on inter-stage dependencies.
+While block-wise T-Recomp successfully prevents dependency conflicts within chunks, inter-chunk dependencies remain critical.
+As illustrated in Fig.
+6(d), block-wise T-Recomp introduces intervals during forward computation.
+As pipeline stage (P) increases, these intervals accumulate—progressively shortening the time between the completion of chunk 1’s forward pass and the initiation of chunk 2’s forward pass, eventually triggering inter-chunk dependency conflicts.
+Inter-chunk dependency conflicts can be effectively resolved.
+When P increases (e.g., $P\geq 8$), delaying chunk 2’s forward pass effectively prevents inter-chunk dependency conflicts.
+Additionally, we demonstrate in Appendix B that for $P\leq 40$, delaying the computation of chunk 2 by just one round is sufficient to avoid dependency conflicts between chunks—a parallelism level that significantly exceeds typical configurations.
+Notably, chunk 2’s forward pass starts later but completes its backward pass earlier than chunk 1, ensuring no impact on total runtime.
+Benefits of T-Recomp
+T-recomp demonstrates better recomputation efficiency.
+We can see that block-wise recomputation introduces some intervals during forward computation, resulting in a slight increase in extrinsic temporal overhead.
+However, because recomputation proportionally extends the execution cycles of workloads in pipeline parallelism, it does not increase storage requirements for deep-layer activations under T-Recomp.
+While this approach requires larger buffer compared to layer-grouped recomputation, it does not significantly impact overall activation storage efficiency.
+These findings are rigorously analyzed in Appendix C.
+In total, the combined storage footprint of activations and buffers is approximately 25% $m_{a}$only one-fourth of the activation storage required by standard 1F1B and a 2× reduction compared to 1F1B+50% recomputation.
+This demonstrates T-Pipe+T-Recomp’s superior memory efficiency.
+T-Recomp demonstrates broader applicability beyond T-Pipe.
+Since its block-wise recomputation design preserves dependency relationships intact, any pipeline parallelism approach exhibiting intra-stage temporal locality variations can effectively adopt T-Recomp.
+Fig.
+6(e) illustrates this generalization through T-Recomp-enhanced Interleave-1F1B scheduling: after recomputing shallow-layer activations with poor temporal locality, the system retains only $\sim$50% $m_{a}$.
+This enables Interleave-1F1B+T-Recomp to match the memory footprint of 1F1B+50% standard recomputation while halving the bubble compared to 1F1B+50% standard recomputation.
+5. TEMPORAL-LOCALITY-AWARE SCHEDULE FOR MODEL STATE
+T-Offload Scheduling Principle
+As outlined in Motivation 1, model states exhibit temporal locality variations across layers similar to activations, enabling prioritized eviction of poor locality model states from HBM.
+Unlike activations that can be recomputed, model states require offloading to DRAM when evicted, an approach we term T-Offload.
+But offloading model states only happens once per mini-batch, resulting in significantly lower data movement volume.
+Quantitative analysis using Llama3’s technical report data (LLAMA3, ) reveals this disparity: for a batch size of 16M tokens, model state size of Llama3-405B is approximately three orders of magnitude smaller than its activation memory footprint.
+Hardware-agnostic T-Offload Solution by Co-design with T-Pipe
+Figure 7. (a)T-Offload in form of overlapping. (b)hardware-agnostic T-Offload in T-Pipe.
+However, implementing T-Offload presents varying challenges across different pipeline parallelism (PP) strategies.
+In existing PP, the communication required for T-Offload overlaps with warm-up/cooldown computations, introducing non-trivial overhead.
 As shown in Fig.
-5(a), with a chunk size of 2, Chronos-Pipe’s scheduling introduces an interval only between the two chunks in both the forward and backward passes.
-Theoretical Analysis
-A further theoretical analysis of Chronos-Pipe’s impact on activation storage requirements and P2P communication is provided.
-For simplicity, we assume $T_{bwd}=2T_{fwd}$ and set $T_{c}$ for a single P2P communication to zero unless stated otherwise.
-With these assumptions, basic unit of time $T_{unit}$ in the pipeline scheduling diagram (e.g., Fig.
-5(a)) is $\frac{T_{fwd}}{2p}$, and the activation storage for each task is $\frac{m_{a}}{2p}$.
-In Chronos-Pipe, activation storage can be estimated by analyzing the activation lifespan, approximately 75% of $m_{a}$.
+7(a), the communication tasks consume a portion of Streaming Multiprocessor (SM) resources on GPUs, which not only increases programming complexity but also degrades computational performance.
+Specifically, competition for SM resources slows down the execution of overlapping compute blocks, and this slowdown introduces additional bubbles during the warm-up and cooldown phases.
+These issues become pronounced in hybrid DP-PP-TP scenarios.
+While dedicated asynchronous communication engines could potentially address these issues, such solutions remain hardware-specific and introduce their own programming complexity.
+In contrast, T-Pipe offers a hardware-agnostic and simple integration solution.
+As illustrated in Fig.
+7(b), T-Pipe naturally introduce bubbles that can be utilized by T-Offload.
+This phenomenon stems from T-Pipe’s core optimization for activation temporal locality: by eliminating intervals between pipeline tasks to maximize locality, the inherent dependencies between chunks automatically generate these bubbles during warm-up and cooldown phases.
+These bubbles enable straightforward T-Offload implementation during idle phases, as shown in Fig.
+7(b).
+During the cooldown phase, we schedule offloading chunk 2’s gradients and performing its optimizer updates between chunk 2’s backward pass completion and chunk 1’s backward pass initiation.
+Quantized weight reloads are allocated to the idle time between two chunk forward computations in the next mini-batch’s warm-up phase.
+This approach preserves throughput with simplicity, avoiding the programming overhead and resource contention inherent to computation-communication overlap—a key advantage over complex overlap-dependent methods.
+Notably, T-Pipe’s bubble exhibits linear growth with pipeline stage (P), confirming that this hardware-friendly T-Offload approach retains scalability.
+Our rigorous analysis in Appendix  D demonstrates that the solution maintains robust scaling with PP,DP and sequence length.
+Furthermore, the optimizer step involves synchronized operations, such as gradient norm calculations, which aggregate all gradients.
+Since T-Offload splits optimizer updates between the CPU and GPU, synchronization overhead must be addressed.
+To mitigate this, a post-optimizer update validation strategy, similar to (ZeroBubble, ), can be employed to manage these challenges effectively.
+Compatibility of T-Offload
+T-Offload operates independently of scheduling strategies by utilizing T-Pipe’s intrinsic idle-phase bubbles for offloading operations.
+This enables native support for tensor, context, and data parallelism, ensuring broad applicability across distributed training environments.
+T-Offload seamlessly integrates with CPU activation Offloading.
+It offloads model states during bubbles, leveraging PCIe upload bandwidth during warm-up and offload bandwidth during cool-down.
+This strategy complements activation Offloading by maximizing PCIe bidirectional bandwidth utilization without contention.
+Therefore, T-Offloading effectively utilizes the temporal locality of weights while also maximizing the use of the prevalent ’bubbles’ during pipeline parallelism’s warm-up and cool-down phases.
+Consequently, it can seamlessly integrate with numerous scheduling optimization strategies.
+Figure 8.
+End-to-End Evaluation on TPipe.
+Input of ColumnLinear is not evenly divided along the TP dimension for better performance.
+(a) shows the memory saving of different schedules on a 48-layer(LLAMA 42B) model at ($P$, TP)=(8,8), global batch=128, micro batch=2, seq len=4K.
+*: R=66.7% in interleave 1F1B, R=50% in 1F1B and T-Recomp is used in T-Pipe.
+(b) illustrates the maximum trainable LLAMA-based model size of different schedules at ($P$, TP)=(8,8), global batch=128, micro batch=2, seq len=4K.
+*: compared with 1F1B.
+6. EXPERIMENTAL RESULTS
+This section addresses three critical research questions:
+$\bullet$ How does TPipe perform in terms of memory efficiency and throughput (Section  6.2)?
+$\bullet$ What are TPipe’s advantages and limitations against other memory-efficient pipeline parallelism approaches (Section  6.3)?
+$\bullet$ What potential gains could be achieved through finer-grained TPipe optimizations (Section  6.4)?
+6.1. Methodology
+Setup.
+Our experiments were carried out by default on a cluster comprising up to 64 GPUs, organized across 8 nodes.
+Nodes are interconnected with four 200 Gbps NICs.
+Each node is equipped with 8 GPUs and 2 CPUs(Intel Xeon Platinum 8480+) and 2TB DRAM, 15.36TB NVME.
+Each GPU features 32GB of HBM and connects to the CPU through PCIe5 x8(32GB/s).
+To thoroughly evaluate the benefits of temporal locality optimization, our experiments require extensive configuration tuning for design space exploration.
+To accelerate this process, we have augmented our experimental validation with large-scale A100 cluster simulations.
+Each GPU in the A100 cluster is equipped with 80GB of HBM memory, with 300GB/s NVLink intra-node connections and 100Gbps inter-node scale-out bandwidth per GPU.
+For performance evaluation, we profile operator execution times while accounting for collective communication overhead, deriving pipeline stage durations and quantify pipeline bubbles and P2P communication using established methods  (SimAI, ).
+Memory requirements become analytically tractable once scheduling strategies are fixed.
+This simulation framework enables precise comparison between TPipe and other SOTA solutions.
+Workloads.
+We evaluate TPipe on two representative models—LLAMA2-70B (LLAMA2, ) and GPT3-175B (GPT3, ).
+Across different experiments, the sequence length varies and model size is adjusted by modifying the number of model layers.
+Baselines.
+TPipe is benchmarked against 1F1B and 1F1B-I(Interleave-1F1B) to quantify the benefits of the temporal locality.
+Recomputation-enabled variants are labeled as ”strategy+R=X%”, e.g., 1F1B+R=50% denotes 50% recomputation.
+Additionally, through simulations, we comprehensively compare TPipe with other memory-efficient pipeline parallelism strategies, including AdaPipe (which employs stage-aware task scheduling) (AdaPipe, ) as well as V-Min and V-Half (which utilize locality-balanced task scheduling) (CtrlMemPP, ).
+When analyzing AdaPipe, we set the same storage budget for AdaPipe and T-Pipe + T-Recomp and compared the differences in their MFU (Model FLOPs Utilization) to conduct a more comprehensive and fair comparison.
+For the following experiments, all evaluations were performed on real GPU clusters unless otherwise indicated.
+Moreover, operator-level recomputation (MegatronKwai, ) and FlashAttention (Flashattention, ) are enabled by default.
+6.2. End-to-End Evaluation
+Fig.
+8(a) demonstrates TPipe’s advantage in HBM demand reduction and throughput preservation.
+Under PP8_TP8, TPipe trains a 48-layer (LLAMA 42B) model where Interleave-1F1B/1F1B trigger OOM.
+T-Recomp slashes activation memory from 16.61GB to 6.39GB, outperforming 1F1B+R=50% (10.65GB activation usage is needed).
+With T-Offload, TPipe ALL achieves 97.58% throughput of 1F1B+R=50% while using only 63.35% of activation and model state memory.
+Moreover, fig.
+8(b) quantifies TPipe’s memory efficiency under PP8_TP8 (32GB HBM): While 1F1B supports only 40 layers (requiring 50% recomputation for 64 layers), T-Pipe trains 48-layer models natively, extendable to 80 layers via T-Recomp and 96 layers (LLAMA 84B) with further T-Offload integration—all while keeping the max reserved memory at roughly the same level.This achieves 2.4× and 1.5× model scale improvements over 1F1B and 1F1B+R=50% respectively, demonstrating temporal locality’s critical role in PP.
+Figure 9.
+memory and throughput of different schedules under various sequence lengths on a 32-layer LLAMA-based model at (DP, $P$, TP)=(2,4,8), global batch=128, micro batch=2.
+*: R=50%, T-Recomp is used in TPipe.
+**: T-Offload is further adopted.
+Input of ColumnLinear is not evenly divided along the TP dimension for better performance.
+We further study the behavior of different scheduling strategies across various sequence lengths, shown in Fig.
+9.
+Under the DP2 (ZeRO-1)_PP4_TP8 configuration, Interleave-1F1B exhibits the worst memory efficiency, running out of memory (OOM) at a sequence length of 8k, even with R=50%.
+Theoretical analysis confirms smaller PP degrades TPipe’s memory gains—achieving only 12.5% (T-Pipe) and 25% (T-Recomp) activation memory reduction versus 1F1B and its variants, but these savings scale markedly with longer sequences.
+T-Offload further reduces model state memory, enabling TPipe to sustain efficiency as both sequence length and model size increase.
+In terms of throughput, while T-Pipe experiences a relative throughput drop of 6%–9% compared to 1F1B, both T-Recomp, and TPipe ALL maintain throughput comparable to 1F1B+R=50% across all sequence lengths despite minor throughput fluctuations from the Operator Library.
+6.3. Analysis about Memory-efficient PP Solutions
+Figure 10. Pareto Frontier of MFU and Mem among Various Schedule@PP8_TP8 A100 cluster. Global batch=32,micro batch=1, seq len=16K.
+Fig.
+10 presents the MFU and memory characteristics across different scheduling strategies, where memory-efficient pipeline parallelism (PP) approaches dominate the Pareto frontier.
+Compared to memory-optimized V-Min and 1F1B+R=50%, T-Pipe+T-Recomp achieves 19.27%-25.14% and 12.23%-16% higher memory efficiency respectively while maintaining comparable or superior MFU.
+When evaluated against AdaPipe, T-Pipe+T-Recomp replaces complex dynamic compilation and search mechanisms with a simpler static approach, incurring only a 2.59%-2.82% MFU reduction at equivalent memory usage.
+The solution further demonstrates seamless integration with T-Offload, where TPipe ALL delivers 26.98%-32.92% memory efficiency gains over AdaPipe, establishing itself as both simpler and more memory-efficient.
+Under relaxed memory constraints, Interleave-1F1B(1F1B-I) with T-Recomp achieves higher MFU than the basic R=50% scheme while saving approximately 25% activation storage, with both Interleave-1F1B(1F1B-I)+T-Recomp and V-Half exhibiting optimal MFU-memory trade-offs that define the Pareto frontier.
+Figure 11.
+MFU of different PP schedule on 48 layers LLAMA-based model@PP8_TP8 A100 cluster when scaling context length.
+global batch=16, micro batch=1.
+Neither T-Recomp nor R=50% solution incorporates recomputation for attention mechanisms for better MFU.
+Figure 12.
+Memory budget of different PP schedule on 48 layers LLAMA-based model@PP8_TP8 A100 cluster when scaling context length.
+global batch=16, micro batch=1.
+Neither T-Recomp nor R=50% solution incorporates recomputation for attention mechanisms for better MFU.
+Fig.
+11 and  12 present the MFU and memory budget under long-sequence training scenarios, where TPipe demonstrates progressively greater computational and memory advantages as sequence length increases.
+Compared to standard 1F1B, the 1F1B-I+T-Recomp configuration achieves 24.58%-38.41% higher memory efficiency while simultaneously improving MFU by 2.75%-4.94%.
+Although V-Half maintains comparable memory efficiency to 1F1B-I+T-Recomp, their MFU gap widens significantly with longer sequences - reaching a 6.14% difference (47.94% vs 54.08%) at 64K sequence length.
+This divergence stems from V-Half’s partitioned backpropagation (BPA and BPW) which introduces increased pipeline bubbles during long-sequence training.
+For similar reasons, T-Pipe+T-Recomp outperforms V-Min by 6.30% MFU (46.13% vs 39.83%) at 64K while maintaining superior memory efficiency.
+Furthermore, as sequences extend from 16K to 128K, the MFU gap between T-Pipe+T-Recomp and AdaPipe narrows from 0.94% to 0.48% under equivalent memory budgets, demonstrating comparable computational efficiency with significantly simpler implementation.
+Figure 13.
+Iteration time of different PP schedule on GPT3-175B when scaling TP size@PP8_TP(8,16,32) on A100 cluster.
+Global batch=32, micro batch=1, seq len=16K.
+assume TP groups all communicate with high bandwidth NVLink.
+In addition to sequence length scaling, model size scaling is equally critical, requiring pipeline parallelism (PP) scheduling to maintain compatibility with large-scale tensor parallelism (TP) for dense models, as demonstrated in Fig.
+13.
+As TP size increases, the TPipe solution exhibits scaling capabilities comparable to 1F1B and Interleave-1F1B(1F1B-I), while V-Half and V-Min show significant performance degradation.
+Compared to Interleave-1F1B(1F1B-I)+T-Recomp and T-Pipe+T-Recomp, V-Half and V-Min transition from a 7.67%/1.72% relative throughput advantage at TP8 to a 4.78%/10.31% deficit at TP32, primarily due to increased communication overhead.
+The backpropagation (BP) partitioning into BPA and BPW not only prevents overlapping of BPA’s reduce-scatter operations but also requires additional all-gather operations for both activation gradients and activations in BPW.
+These compounded communication costs fundamentally limit their scalability along the TP dimension.
+6.4. Evaluation on Finer-grained Variants of TPipe
+Figure 14. in-depth analysis of T-Pipe/1F1B-I + T-Recomp. Additional buffer required for recomputation is ignored for better illustration.
+Fig.
+14 shows the effectiveness of T-Recomp across different chunk configurations in T-Pipe and Interleave-1F1B(1F1B-I).
+T-Pipe’s locality-aware scheduling inherently achieves lower memory usage than 1F1B even without recomputation.
+When combined with T-Recomp, it further reduces activation storage at all recomputation ratios compared to standard 1F1B recomputation, with finer-grained chunks showing superior optimization due to more precise temporal locality application.
+While standard Interleave-1F1B underperforms 1F1B at all recomputation ratios due to significant extrinsic temporal overhead, T-Recomp fundamentally solves this problem: by discarding chunks with the poorest temporal locality, T-Recomp enables Interleave-1F1B to match standard 1F1B’s recomputation efficiency, significantly improving its memory utilization.
+Figure 15. (a)evaluate the scalability of T-Offload on a 16-layer model at global batch=128, micro batch=2. (b)in-depth analysis of T-Offload
+We further verify the scalability of T-Offload, as illustrated in Fig.
+15(a).
+When training a 16-layer model with a sequence length of 4K on PP4_TP8 setup, T-Pipe achieves only partial overlap (45.45%) between the bubble created during the cooldown phase and the time required for T-Offload.
+However, when either the pipeline stage or the sequence length is doubled, the overlap improves significantly to 94.55% and 100%, respectively.
+Furthermore, T-Offload enables only half of the optimizers to be updated on GPU, leading to a slight decrease in overall training time.
+These results highlight the scalability of T-Offload, demonstrating its increasing effectiveness as the PP stage and sequence length grows.
+Under multi-chunk partitioning, T-Pipe+T-Offload benefits from increased available bubbles.
+Fig.
+15(b) demonstrates this progressive optimization: when sequence length grows from 4K to 8K, chunk 2 shows only marginal improvement in T-Offload efficiency since existing bubbles already suffice for its model state transfer.
+In contrast, chunks 4 exhibit significant gains as sequence length increases to 8K and beyond, since these chunks initially lack sufficient bubbles and require longer sequences to create adequate offloading windows.
+Chunk 3 exhibits intermediate behavior:at 8K sequences it reaches optimal efficiency with only the first chunk’s poorest-temporal-locality model states remaining unoffloaded.
+When sequence length expands to 16K, chunks 2 and 3 already reach their optimization limits, reducing model state storage by 33.33% and 44% respectively, while chunk 4 achieves near-maximal 49.09% reduction.
+7. RELATED WORK
+Hybrid Parallelism.
+Modern distributed training combines data (DP) (ZeRO, ; FSDP, ), pipeline (PP) (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ), tensor (TP) (AccPar, ; 2D-TP, ; 2p5D-TP, ; 3D-TP, ), and context parallelism (CP) (RingAttention, ; Ulysses, ) to scale model size and sequence length.
+DP/PP/TP distribute model states, while CP/SP (SeqParallel, ) handle activations.
+Hybrid parallelism combines these approaches, enabling efficient scaling to models with trillions of parameters.
+PP’s low communication makes it critical for cross-node deployments.
+However, it suffers from bubbles and high activation memory overhead.
+Pipeline Parallelism.
+Most PP research focuses on reducing pipeline bubbles by finer-grained task partitioning  (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ) or multi-dimensional task scheduling  (Chimera, ; Mixpipe, ).
+However, they neglect the significant overhead of activation memory.
+(CtrlMemPP, ) reduces activation lifespans through scheduling but risks introducing more bubbles and prolonging gradient lifespan.
+(AdaPipe, ) addresses activation memory imbalance with asymmetric partitioning but increases programming complexity.
+(BreadthFirstPP, ; ZeroPP, ), achieve ZeRO-2/3-compatibile PP at the cost of higher activation memory.
+(LLAMA3, ) takes advantage of  (BreadthFirstPP, ) and 1F1B for improved ZeRO-2 compatibility but only maintains, rather than reduces, activation memory usage.
+Recomputation and Offloading.
+Memory requirements can also be reduced by trading memory for computation or communication.
+For recomputation, some studies  (SeqParallel, ; MegatronKwai, ; SuperNeuron, ; CapuChin, ) focus on selective recomputation of different operators but fail to address the high memory usage of the projection operator.
+(DTR, ) introduced temporal locality-aware recomputation principle and demonstrated its efficacy in DP, but this principle is still poorly explored in PP.
+Some approaches leverage the characteristics of hybrid parallelism.
+(AdaPipe, ) adapts recomputation to mitigate activation memory imbalance in PP.
+(OverlapRecomp, ) overlaps recomputation with communication to reduce overhead.
+However, these approaches increase programming complexity.
+For offloading,  (Bpipe, ) addresses memory imbalance in PP by offloading activation across GPUs.
+Most studies focus on fully utilizing storage systems.
+For example,  (ZeRO-Offload, ; PatrickStar, ; Mobius, ) explore offloading activation and model states to DRAM, while  (ZeRO-Infinity, ; Angel-ptm, ; Smart-Infinity, ; Fuyou, ) extend this to NVMe.
+Due to limited bandwidth in offloading, these efforts primarily focus on identifying opportunities to overlap offloading with computation and realize near-data processing.
+8. CONCLUSION
+This work introduces TPipe to address the memory capacity limitations in LLM training.
+The core idea behind TPipe is to treat HBM as a fast but limited ”cache,” optimizing and leveraging temporal locality in LLM training for efficient HBM usage.
+T-Pipe enhances activation temporal locality in PP scheduling and, together with T-Recomp and T-Offload, discards data with poor temporal locality from HBM.
+Experiments show that T-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
+At the same time, TPipe has good enough compatibility to make it suitable for large-scale hybrid parallelism training.
+Appendix A Theoretical Analysis of T-Pipe
+Figure 16. Theoretical Analysis of T-Pipe.
+A further theoretical analysis of T-Pipe on activation storage and bubble ratio is provided as follow.
+For simplicity, we assume $T_{bwd}=2T_{fwd}$, basic unit of time $T_{unit}$ in the pipeline scheduling diagram (e.g., Fig.
+16) is $\frac{T_{fwd}}{2p}$, and the activation storage for each task is $\frac{m_{a}}{2p}$.
+In T-Pipe, activation storage can be estimated by analyzing the activation lifespan, approximately 75% of $m_{a}$.
 As shown in Fig.
-5(a), Stage 0 launches chunk 2 of the fourth micro-batch before releasing the activation storage from chunk 2 of the third micro-batch.
+16, Stage 0 launches chunk 2 of the fourth micro-batch before releasing the activation storage from chunk 2 of the third micro-batch.
 Since tasks on a single stage execute in cycles of $6T_{unit}$, the accumulated activation storage for chunk 2 can be represented as $\lceil\frac{T_{life\_chunk2}}{6T_{unit}}\rceil$, where $T_{life\_chunk2}$ denotes the lifespan of chunk 2’s activations, precisely the time from the completion of its forward pass to the beginning of the corresponding backward pass at PP stage 0.
@@
 There is also an extra interval $T_{fwd\_interval}$ between the forward chunks, derived from steps 1, 2, and 3 in Fig.
-5(a) as $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$.
+16 as $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$.
 Similarly, an additional interval $T_{bwd\_interval}$ exists between the backward chunks, derived from steps 4, 5, and 6 in Fig.
-5(a) as $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$.
-Thus, for Stage 0, the lifespans of the activations for chunk 1 and chunk 2, denoted $T_{life\_chunk1}$,$T_{life\_chunk2}$ can be expressed as:
+16 as $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$.
+Thus, for Stage 0, the lifespans of the activations for chunk 1 and chunk 2, denoted $T_{life\_chunk1}$$T_{life\_chunk2}$ can be expressed as:
 (1) $T_{life\_chunk2}=T_{fwd\_chunk2}+T_{bwd\_chunk2}-2T_{unit}$
@@
 $T_{life\_chunk2}+T_{bwd\_interval}+T_{bwd\_chunk1}$
-As a result, the peak activation storage requirements for chunks 1 and 2 are $\lceil\frac{2}{3}+\lceil\frac{(p-3)}{6}\rceil+\lceil\frac{(2p-3)}{6}\rceil+\frac{p}{2}\rceil\frac{m_{a}}{2p}$,$\lceil\frac{(3p-2)}{6}\rceil\frac{m_{a}}{2p}$ respectively.
+As a result, the peak activation storage requirements for chunks 1 and 2 are $\lceil\frac{2}{3}+\lceil\frac{(p-3)}{6}\rceil+\lceil\frac{(2p-3)}{6}\rceil+\frac{p}{2}\rceil\frac{m_{a}}{2p}$$\lceil\frac{(3p-2)}{6}\rceil\frac{m_{a}}{2p}$ respectively.
 When P is large, these values approach $\frac{m_{a}}{2},\frac{m_{a}}{4}$, making the total peak activation storage approximately 75% of $m_{a}$.
-In Chronos-Pipe, when the task is divided into v chunks, the bubble overhead from P2P communication is approximately v times that of the 1F1B schedule.
-The scheduling of Chronos-Pipe with two chunks under the condition $T_{c}\neq 0$ is considered to illustrate this.
-As shown in Fig.
-5(b), the Chronos-Pipe execution time can be divided into three phases: $T_{warmup},T_{steady}$ and $T_{cooldown}$.
-$\bullet$ The warmup phase primarily includes the forward pass of the first micro-batch, calculated as $(2p+2)T_{unit}+(2p+1)T_{c}$.
-$\bullet$ The Steady phase covers the forward pass of chunk 1 for the remaining $(m-2)$ microbatches, totaling $(m-2)$$(6T_{unit}+8T_{c})$.
-$\bullet$ The cooldown phase mainly involves the backward pass of the last micro-batch, calculated as $(4p+4)T_{unit}+(2p+3)T_{c}$.
-Adding the time of these three phases together, the total time is $6(m+p-1)T_{unit}+(4p+8m-12)T_{c}$.
-Exactly, Fig.
-5(b) shows that the warmup and cooldown phases can each remove one $T_{c}$ of bubble overhead, leading to an adjusted total time of $6(m+p-1)T_{unit}+(4p+8(m-2)+2)T_{c}$.
-In comparison, the execution time for 1F1B is derived as $6(m+p-1)T_{unit}+(2p+4(m-2))T_{c}$.
-From this, we can see that Chronos-Pipe with two chunks has a P2P communication overhead slightly greater than twice that of 1F1B.
-This is because P2P communication between Stage 0 and Stage (P-1) appears in the critical path of the Chronos-Pipe schedule during the warmup and cooldown phases.
-This analysis holds for other values of $v$, where the P2P communication bubble overhead will slightly exceed v times that of 1F1B.
-Further analysis shows that the theoretical bubble overhead for Chronos-Pipe with two chunks is:$\frac{6(p-1)T_{unit}+[4p+8(m-2)+2]T_{c}}{6(p-1+m)T_{unit}+[4p+8(m-2)+2]T_{c}}$ whereas the bubble overhead for 1F1B is: $\frac{6(p-1)T_{unit}+[2p+4(m-2)]T_{c}}{6(p-1+m)T_{unit}+[2p+4(m-2)]T_{c}}$Set $T_{c}=0$, the bubble overhead for Chronos-Pipe matches that of 1F1B.
-When $T_{c}=0.05T_{unit},m=128,p=4$, the theoretical bubble overheads are 8.27% for Chronos-Pipe and 5.37% for 1F1B.
-Due to the increased P2P communication in Chronos-Pipe, the Model Flop Utilization (MFU) is slightly reduced.
-Implementation of Chronos-Pipe
-In Chronos-Pipe’s implementation, partial asynchronous communication is introduced to reduce bubble size.
-In mainstream frameworks like DeepSpeed and Megatron, the 1F1B configuration uses synchronous P2P communication.
-However, in Chronos-Pipe, there can be significant intervals between chunks during forward and backward passes.
-If synchronous P2P communication is used in these intervals, much of the time would be spent waiting for communication to complete.
-To address this, Chronos-Pipe modifies P2P communication between chunks to be asynchronous, thereby achieving higher throughput.
-As illustrated in Fig.
-5(b), after Stage 0 completes the backward computation for chunk 2, it can proceed with subsequent tasks until just before Stage (P - 1) begins the corresponding backward pass for chunk 1, at which point the P2P communication is completed.
-Other P2P communications remain synchronous, which is consistent with mainstream frameworks.
-Figure 6. (a) build Chronos-Recomp based on Chronos-Pipe. (b) Chronos-Recomp Schedule with PP stage is set to 8.
-4.2. Chronos-Recomp: Chronos-aware Recomputation
-Recompute Shallow Layer First!
-Chronos-Pipe aims to shorten the lifespan of activations in shallow layers by reducing the interval.
-Yet, it cannot fully eliminate the intrinsic differences in temporal locality across layers within DNN.
-We can further mitigate this issue by discarding data with low temporal locality from the fast but limited HBM.
-This approach suggests prioritizing the recomputation of shallower layers—an approach we refer to as Chronos-Recomp—when recomputation is considered.
-Chronos-Recomp can be seamlessly built based on Chronos-Pipe.
-We can reserve time for recomputation just before the backward computation block of the shallow layers, as shown in Fig.
-6(a).
-This approach introduces regular intervals during the forward pass without changing the overall pipeline scheduling.
-Therefore, Chronos-Recomp can also support partial recomputation by controlling the length of the interval.
-As discussed in the theoretical analysis of Chronos-Pipe, the intervals between Chronos-Pipe chunks depend on the value of P.
-As P increases, dependency conflicts may arise in this approach.
-However, these conflicts can be avoided by delaying the forward computation launch of chunk 2, as illustrated in Fig.
-6(b) for Chronos-Recomp scheduling with P=8.
-This adjustment does not affect the bubble ratio and memory usage, which will be explained in the following theoretical analysis of Chronos-Recomp.
-Theoretical Analysis
+In addition, T-Pipe exhibits a bubble ratio similar to 1F1B. As depicted in Fig.  16, its execution comprises three phases: $T_{warmup},T_{steady}$ and $T_{cooldown}$.
+$\bullet$ The warmup phase includes the forward pass of the first micro-batch, calculated as $2pT_{unit}$.
+$\bullet$ The Steady phase covers the forward pass of chunk 1 for the remaining (m1)(m-1) microbatches, totaling $6(m-1)T_{unit}$.
+$\bullet$ The cooldown phase mainly involves the backward pass of the last micro-batch, calculated as $4pT_{unit}$.
+The total execution time aggregates three phases into $6(m+p-1)T_{unit}$, with bubble time accounting for $6(p-1)T_{unit}$.
+Consequently, the bubble ratio is $\frac{p-1}{m+p-1}$—identical to 1F1B when P2P communication costs is ignored.
+In fact, the partition of workload into two chunks requires twice the P2P communication rounds compared to 1F1B.
+Due to increased P2P communication in T-Pipe, the Model Flop Utilization is slightly reduced.
+Appendix B Formalization on Delay Rounds at T-Recomp
+In T-Pipe scheduling, an additional interval time $T_{fwd\_interval}$ exists between forward chunks, derived from steps 1, 2, and 3 in Fig.
+16, resulting in $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$.
+In T-Recomp’s shallow-layer full recomputation, the execution times for steps 1, 2, and 3 respectively change $T_{unit},\lceil\frac{(p-3)}{6}\rceil T_{unit}$, and $\lceil\frac{(p-1)}{2}\rceil T_{unit}$.
+Therefore, $\Delta T_{fwd\_interval}$ becomes $(\lceil\frac{(p-1)}{2}\rceil-\lceil\frac{(p-3)}{6}\rceil-1)T_{unit}(P\geq 3)$.
+Based on this, We can formalize the number of rounds k required to delay the launch of chunk 2 as an optimization problem for $P\geq 3$:
+(3) $Min$ $\quad k$
+$s.t.\quad T_{fwd\_interval}$- $\Delta T_{fwd\_interval}+7kT_{unit}\geq 0$
+$k$ $\in N$
+$p$ $\geq 3$
+Therefore, when $8\leq p\leq 40$, we have $k=1$, meaning that chunk 2 needs to be delayed by one round.
+Appendix C Theoretical Analysis of T-Recomp
 We analyze activation storage requirements using full recomputation of shallow layers as an example.
@@
 Observing micro-batch 2 in Fig.
-6(b), the total intervals introduced during the forward execution of chunk 2 is $\lceil\frac{(p-1)}{2}\rceil T_{unit}$.
+6(d), the total intervals introduced during the forward execution of chunk 2 is $\lceil\frac{(p-1)}{2}\rceil T_{unit}$.
 Therefore, for Stage 0, the lifespan of chunk 2’s activations is given by:
 (4) $T_{life\_chunk2}=(3p+\lceil\frac{(p-1)}{2}\rceil-2)T_{unit}$
-Here, $(3p-2)T_{unit}$ represents the ideal time from the start of chunk 2’s forward computation to the arrival of the backward computation, assuming no interval is present.
+Here, $(3p-2)T_{unit}$ represents the intrinsic temporal overhead of chunk 2.
 Moreover, since tasks in pipeline parallelism repeat with a period of $7T_{unit}$, the required number of activation storage blocks for chunk 2 is $\lceil\frac{T_{life\_chunk2}}{7T_{unit}}\rceil$, which simplifies to $\lfloor\frac{p}{2}\rfloor$.
-Thus, with Chronos-Recomp’s full recomputation for chunk 1, the remaining activation storage is only 25% of $m_{a}$.
-By comparison, standard recomputation in 1F1B with the same computation budget requires 50% of $m_{a}$ for storage, making Chronos-Recomp 1.5 times more storage-efficient than standard recomputation.
-Under these conditions, we can also evaluate the total execution time, where Chronos-Recomp is similar to 1F1B with recomputation, as shown in Fig.
-6(b).
-The warm-up phase, $T_{warmup}$, represents the time required to complete the forward computation of chunk 1, taking $2pT_{unit}$.
-The steady phase, $T_{steady}$, involves the forward computation of chunk 2 for the remaining $(m-1)$ microbatches, consuming $7(m-1)T_{unit}$.
-During the cool-down phase, the backward computation of chunk 1 for the final micro-batch requires $4pT_{unit}$.
-This results in a total execution time of $[6p+7(m-1)]T_{unit}$.
+Taking into account that the recomputation buffer requires an additional block of space for regenerated activation values, we need $\lfloor\frac{p}{2}\rfloor+1$ blocks for activation.
+Thus, with T-Recomp’s full recomputation for chunk 1, the activation storage is only 25% of $m_{a}$.
+By comparison, standard recomputation in 1F1B with the same computation budget requires 50% of $m_{a}$ for storage, making T-Recomp 1.5 times more storage-efficient than standard recomputation.
+Execution time of T-Pipe+T-Recomp is similar to 1F1B with recomputation, as shown in Fig.
+6(d).
+The warm-up phase/cool-down phase take $2pT_{unit}$$4pT_{unit}$ respectively.
+The steady phase involves the forward computation of chunk 2 for the remaining (m1)(m-1) microbatches, consuming $7(m-1)T_{unit}$.
+This results in a total execution time of $6p+7(m-1)]T_{unit}$.
 By comparison, the execution time for 1F1B with 50% recomputation is $7(m-1+p)T_{unit}$.
-Notably, Chronos-Recomp’s forward computation allows for overlap with half of the P2P communications.
-In practical scenarios, where $T_{bwd}\leq 2T_{fwd}$, the P2P communication of gradient in the backward pass of chunk 2 can overlap with recomputation.
-Consequently, the impact of P2P communication on Chronos-Recomp and 1F1B is similar.
-As P increases, dependency conflicts can arise within the Chronos-Recomp scheduling scheme.
-We’ve analyzed this situation in detail.
-When P increases (e.g., $P\geq 8$), delaying the launch of the forward computation for chunk 2 effectively prevents dependency conflicts between the forward computations of the two chunks.
-Additionally, we demonstrate in Appendix A that for $P\leq 40$, delaying the computation of chunk 2 by just one round is sufficient to avoid forward dependency conflicts between chunks—a parallelism level that significantly exceeds typical configurations in mainstream settings.
-Notably, chunk 2’s forward execution begins later than chunk 1’s, while its backward pass completes earlier.
-This scheduling adjustment, therefore, does not impact the critical path and does not affect the overall execution time.
-Moreover, the peak activation storage analysis we conducted does not impose any limitations on the value of P, making the conclusions widely applicable across various P values.
-4.3. Compatible with ZeRo-2 Based Pipeline Schedule
-Figure 7. build Zero-2-compatible schedule based on Chronos-Pipe
-Chronos-Pipe can be further adapted to support a ZeRO-2-compatible PP.
-As illustrated in Fig.
-7, the modified Chronos-Pipe repeatedly launches the workload for the current chunk’s forward computation.
-This increases the overlap time available for ZeRO-2’s communication between DP ranks.
-Additionally, this adjustment has minimal impact on activation storage, enabling the launch of additional workloads of the same chunk to reduce communication demands between DP ranks.
-Furthermore, When recomputation needs to be considered, such as long sequence training, Chronos-Recomp can offer higher recomputation efficiency for this ZeRO-2-compatible PP scheduling approach.
-5. Chronos-aware SCHEDULE FOR MODEL STATE MEMORY SAVINGS
-Temporal locality can be applied not only to activation memory usage optimization but also to model state store optimization.
-5.1. Chronos-Offload: Chronos-aware Offload
-Chronos-Offload Scheduling Principle
-The central idea behind Chronos-Offload is to offload model states with poor temporal locality to the CPU.
-In DNN, weights of deeper layers complete gradient computation first, allowing them to be updated earlier.
-However, these weights are the last to be used in the forward pass, giving them the poorest temporal locality.
-In the two-chunk Chronos-Pipe design, this is reflected by chunk 2 starting its forward computation later than chunk 1 but finishing its backward pass earlier.
-As a result, chunk 2’s weight updates can be offloaded to the CPU, while chunk 1’s weight updates remain on the GPU.
-This approach allows us to overlap chunk 2’s weight update process with the idle time between the end of its backward pass in the current mini-batch and the start of its forward pass in the next mini-batch.
-This approach aligns well with the Optimizer Step phase’s specific computation and storage characteristics.
-In this phase, gradients are used to update first- and second-order momenta and the master weight (32-bit).
-The updated master weight is then quantized to produce a 16-bit quantized weight.
-Since this process involves only element-wise operations, it doesn’t leverage the high computational power of Tensor Cores, resulting in inefficient GPU utilization.
-On the other hand, server CPUs, with their SIMD processing capabilities, are well-suited to the computational demands of the Optimizer Step.
-By offloading weight updates to the CPU, the first- and second-order momenta and the master weight can be kept in CPU-side DRAM, with only the quantized weight and gradients stored on the GPU.
-This reduces the GPU’s storage requirements to roughly one-third of the total Model State.
-Through Chronos-Offload, we introduce temporal locality into the Model State Offload strategy and make it more suitable with PP, enabling complementary use of the CPU and GPU’s respective computing and memory strengths while minimizing impact on GPU computational efficiency.
-Implement of Chronos-Offload
-Chronos-Pipe offers seamless compatibility with Chronos-Offload in pipeline parallelism compared to other methods.
-In existing approaches, the cooldown phase immediately transitions from chunk 2’s backward pass to chunk 1’s backward pass, eliminating scheduling bubbles.
-A similar issue arises in the warm-up phase, as seen in the interleaved 1F1B structure (Fig.
-2(a)).
-Supporting Chronos-Offload in this setup requires modifying the PyTorch backend to asynchronously manage chunk 2’s weight updates on the CPU, including gradient offloading, optimizer updates, and updated weight uploads.
-These intrusive changes increase programming complexity and risk resource conflicts, reducing compatibility with other scheduling strategies.
-In contrast, Chronos-Pipe naturally supports Chronos-Offload by introducing bubbles between tasks.
-During the cooldown phase, bubbles exist between the end of chunk 2’s backward pass and the start of chunk 1’s, with a similar effect in the warm-up phase.
-These bubbles provide ideal opportunities for Chronos-Offload without interference.
-This behavior stems from Chronos-Pipe’s optimization of temporal activation locality, minimizing intervals between pipeline tasks while naturally creating sufficient bubbles during warm-up and cooldown phases.
-Figure 8. seamless compatibility with Chronos-Offload in Chronos-Pipe
-Chronos-Pipe utilizes these bubbles to enable smooth integration with Chronos-Offload, as shown in Fig.
-8.
-During the cooldown phase, gradients of chunk 2 are offloaded, and the corresponding optimizer updates are performed in the bubbles between chunk 2’s backward pass and chunk 1’s.
-Similarly, the upload of quantized weights is scheduled during bubbles between forward passes of the two chunks in the next minibatch’s warm-up phase.
-While this approach doesn’t fully optimize Chronos-Offload’s efficiency, we demonstrate that its effectiveness improves with higher levels of pipeline parallelism (PP), data parallelism (DP), and longer sequence lengths.
-Furthermore, the optimizer step involves synchronized operations, such as gradient norm calculations, which aggregate all gradients.
-Since Chronos-Offload splits optimizer updates between the CPU and GPU, synchronization overhead must be addressed.
-To mitigate this, a post-optimizer update validation strategy, similar to (ZeroBubble, ), can be employed to manage these challenges effectively.
-Scalability with Increasing DP, PP, and Sequence Length
-Two conditions must be met to complete the Chronos-Offload process within the available ’bubbles.’ These conditions are more easily satisfied as the degree of PP and sequence length increase.
-The first condition requires the gradient offload and the CPU optimizer updates to be completed within the cooldown phase’s ’bubbles.’ In Chronos-Pipe, the interval between backward computations of chunks is $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$, which results in a delay of $\lceil\frac{(2p-3)}{6}\rceil$ rounds for chunk 1 relative to chunk 2.
-This aligns the backward computation of the first micro-batch for chunk 1 with the (1+$\lceil\frac{(2p-3)}{6}\rceil$) th one of chunk 2.
-Thus, we define:
+Notably, T-Recomp achieves 50% P2P communication overlap during forward computation, resulting in similar P2P communication impacts between T-Pipe+T-Recomp and 1F1B.
+Appendix D Scalability of T-Offload with Increasing DP, PP, and Sequence Length
+[…]
 (5) $T_{avaliable\_offload}=(p-\lceil\frac{(2p-3)}{6}\rceil-1)\frac{T_{bwd}}{2p}$
@@
 The ratio $\frac{T_{step}}{T_{bwd}}$ is constant for a given model and is influenced by factors such as sequence length, batch size, offload communication bandwidth, and CPU processing speed.
-Longer sequences and larger batch sizes further decrease $\frac{T_{step}}{T_{bwd}}$, linearly increasing the idle time available for Chronos-Offload, even without considering the attention operator.
+Longer sequences and larger batch sizes further decrease $\frac{T_{step}}{T_{bwd}}$, increasing the bubble available for T-Offload.
 The second requirement is that quantized weight uploads must be completed within the ’bubbles’ of the warm-up phase.
@@
 Similar to the first condition, this requirement becomes easier to meet with increased pipeline stages p and longer sequence lengths.
-Thus, Chronos-Offload exhibits strong scalability with increasing degrees of pipeline parallelism (PP) and sequence lengths, which is suitable for future development trends.
-Additionally, Chronos-Offload scales effectively with increased Data Parallelism (DP).
+Thus, T-Offload exhibits strong scalability with increasing degrees of pipeline parallelism (PP) and sequence lengths, which is suitable for future development trends.
+Additionally, T-Offload scales effectively with increased Data Parallelism (DP).
 As the level of DP increases, the amount of weights each DP rank needs to update decreases, which in turn reduces the time required for gradient offloading and optimizer computations.
-5.2. Compatible with Other Schedule
-Chronos-Offload is designed to be compatible with a variety of scheduling strategies.
-Our approach exclusively leverages the ’bubbles’ created within Chronos-Pipe, eliminating competition over communication bandwidth and CPU with other scheduling approaches.
-As a result, it seamlessly supports many parallelism strategies, including tensor parallelism, context parallelism, and data parallelism.
-This flexibility makes it widely applicable across diverse distributed parallelism training frameworks.
-Chronos-Offload is also fully compatible with the CPU Activation Offloading scheme.
-It performs Model State Offloading in the interval between the completion of chunk 2’s backward computation and the onset of the forward computation for the next mini-batch of chunk 2.
-This process exclusively utilizes the PCIe upload bandwidth during the warm-up phase and the PCIe offload bandwidth during the cool-down phase.
-This approach does not interfere with CPU Activation Offloading; rather, it enhances the efficient utilization of the PCIe bidirectional bandwidth.
-Therefore, Chronos-Offloading effectively utilizes the temporal locality of weights while also maximizing the use of the prevalent ’bubbles’ during pipeline parallelism’s warm-up and cool-down phases.
-Consequently, it can seamlessly integrate with numerous scheduling optimization strategies.
-6. EXPERIMENTAL RESULTS
-In this section, we seek to answer the following questions:
-$\bullet$ How well does ChronosPipe perform in memory and throughput?How does it compare with other recomputation-based solutions?
-$\bullet$ How does ChronosPipe perform across varying sequence lengths and model architectures (input embedding, attention, activation function)?
-$\bullet$
-What is the actual impact of P2P communication on ChronosPipe? Can the scalability of Chronos-Offload truly maintain performance without degradation?
-$\bullet$ What potential benefits could be achieved by further leveraging temporal locality in ChronosPipe?
-Figure 9.
-End-to-End Evaluation on ChronosPipe.
-(a) shows the memory saving of different schedules on a 48-layer model at (PP, TP)=(8,8), global batch=128, micro batch=2, seq len=4K.
-*: R=66.7% in interleave 1F1B, R=50% in 1F1B and Chronos-Recomp is used in Chronos-Pipe.
-(b) illustrates the maximum trainable model size of different schedules at (PP, TP)=(8,8), global batch=128, micro batch=2, seq len=4K.
-*: compared with 1F1B.
-6.1. Methodology
-Setup.
-Our experiments were carried out on a cluster comprising up to 64 GPUs, organized across 8 nodes.
-Nodes are interconnected with four 200 Gbps NICs.
-Each node is equipped with 8 GPUs and 2 CPUs(Intel Xeon Platinum 8480+) and 2TB DRAM, 15.36TB NVME.
-Each GPU features 32GB of HBM and connects to the CPU through PCIe5 x8(32GB/s), with high-bandwidth interconnects linking GPUs within a node.
-Workloads.
-For these experiments, we utilized a model architecture similar to LLAMA2-70B (LLAMA2, ), consisting of 80 transformer layers and incorporating the GQA mechanism.
-We varied the number of transformer layers in the model to adjust the size for different experiments.
-We further analyze the effectiveness of ChronosPipe on other models through theoretical analysis.
-Qwen2.5-32B (Qwen, ), PaLM-62B (PaLM, ), and OPT-66B (OPT, ) are selected as they have distinct vocabulary sizes, attention mechanism configurations, and activation functions, being typical enough for current dense models.
-Baselines.
-We compare ChronosPipe against 1F1B and interleave-1F1B to highlight the benefits of introducing temporal locality into pipeline parallelism.
-All scheduling strategies were implemented within a framework similar to Megatron-LM to ensure a fair comparison.
-In ChronosPipe’s implementation, only essential asynchronous communications were employed, while all other P2P communications followed the synchronous approach used in the baselines.
-When recomputation is included, we denote it using ”scheduling strategy + recomputation ratio.” For instance, $1F1B+R=50\%$ represents the 1F1B with a 50% recomputation ratio.
-Additionally, through theoretical analysis, we further compare our approach with the recomputation strategies of Megatron-Kwai (MegatronKwai, ) and AdaPipe (AdaPipe, ).
-Conditions are aligned with AdaPipe’s publicly - available data for a fair comparison.
-Unless otherwise specified, the input of ColumnLinear is not evenly divided along the TP dimension, consistent with Megatron’s settings.
-Meanwhile, Operator-level recomputation (RMSNorm, activation function) and FlashAttention (Flashattention, ) are enabled by default.
-In theoretical calculations, $T_{c}=0.104T_{unit}$ is set.
-Theoretical analysis.
-Based on the following two reasons, we assert that theoretical analysis demonstrates sufficient credibility in computational and storage evaluation.
-First, the performance of PP can be effectively modeled.
-By precisely quantifying the bubble ratio and recomputation proportion, accurate performance estimation becomes achievable—an approach validated as feasible in  (PipeFisher, ; Chimera, ).
-Regarding storage, Activation and Model State dominate storage consumption, enabling theoretical analysis to effectively identify critical storage constraints.
-6.2. End-to-End Evaluation
-Fig.
-9(a) illustrates that ChronosPipe significantly reduces HBM memory requirements while maintaining comparable throughput.
-Under the PP8_TP8 configuration, ChronosPipe’s optimized scheduling enables training a 48-layer (i.e., 42B) model, while Interleave-1F1B and 1F1B run into out-of-memory (OOM) issues.
-Chronos-Recomp further demonstrates improved recomputation efficiency, reducing activation memory from 17.23 GB to 6.63 GB.
-In contrast, under the same recomputation budget (R=50%), the 1F1B method still occupies 10.6 GB for activations.
-With Chronos-Offload, ChronosPipe ALL reduces the total memory footprint for activations and model states to 64.6% of what is used in the $1F1B+R=50\%$ setup while achieving 97.58% of its throughput.
-Fig.
-9(b) explicitly evaluates the effect of ChronosPipe in terms of the maximum size of the trainable model.
-With limited 32GB HBM, 1F1B can train only a 40-layer model under the PP8_TP8 configuration, while training a 64-layer model requires a 50% recomputation ratio.
-In contrast, ChronosPipe can train a 48-layer model, and with the Chronos-Recomp strategy, this can be extended to 80 layers.
-Further leveraging Chronos-Offload enables training even a 96-layer model, all while keeping the max reserved memory at roughly the same level.
-As a result, ChronosPipe increases the scale of trainable models to 2.4 times the 1F1B, and 1.5 times that of the 1F1B+R=50% configuration.
-Note that this only shows the advantage of introducing temporal locality into PP.
-ChronosPipe can further integrate other compatible optimization strategies to optimize HBM usage, which is not reflected in this experiment.
-Figure 10.
-Estimated memory saving for different recomputation-based solutions on GPT3-175B at (PP,TP)=(8,8), global batch=32, micro batch=1, seq len=16K.
-sequence parallelism (SP) (SeqParallel, ) is used in all and operator-level recomputation(RMSNorm, activation function) is not used in ChronosPipe.
-*:1F1B with only Operator-level recomputation solution is used.
-activation-offloading is not under consideration due to it suitable for other solutions.
-Fig.
-10 compares ChronosPipe with other recomputation solutions.
-Megatron-Kwai’s Operator-aware recomputation reduces storage by 1.27x with negligible performance loss.
-Applying Operator-aware recomputation to remaining activations shows diminishing returns, but co-designing PP with recomputation can enhance its benefits.
-AdaPipe adopts Stage-aware task and recomputation budget allocation across pipeline stages, achieving 1.76x storage savings and 1.26x performance improvement over the 1F1B+R = 100% baseline.
-Without Operator-aware recomputation on activations of deeper layers, Chronos-Pipe+Chronos-Recomp attains 1.72x storage savings and 1.17x performance gain relative to the same baseline.
-ChronosPipe ALL further achieves 2.22x storage savings, highlighting the efficiency of Locality-aware design.
-While Chronos-Pipe+Chronos-Recomp exhibits moderately reduced throughput relative to AdaPipe, it provides a general and simple solution, and enhance adaptability to a variety of changes including model scale, structures, and parallelism configurations.
-In contrast, AdaPipe necessitates sophisticated search mechanisms and programmatic overhead to implement stage-aware task and recomputation budget allocation, which is a hard integer partitioning problem.
-Maintaining workload equilibrium across varying constraints mentioned while preserving throughput remains questionable.
-Moreover, these approaches embody distinct design philosophies, revealing synergistic potential for future integration.
-Figure 11.
-memory and throughput of different schedules under various sequence lengths on a 32-layer model at (DP, PP, TP)=(2,4,8), global batch=128, micro batch=2.
-*: R=50%, Chronos-Recomp is used in ChronosPipe.
-**: Chronos-Offload is further adopted.
-We further study the behavior of different scheduling strategies across various sequence lengths, shown in Fig.
-11.
-Under the DP2 (ZeRO-1)_PP4_TP8 configuration, Interleave-1F1B exhibits the worst memory efficiency, running out of memory (OOM) at a sequence length of 8k, even with R=50%.
-Theoretical analysis suggests that smaller PP reduces ChronosPipe’s memory efficiency, resulting in Chronos-Pipe and Chronos-Recomp saving only 12.5% and 25% of activation respectively compared to 1F1B and its recomputation variants in this setup.
-However, as sequence lengths increase, ChronosPipe’s memory savings become increasingly pronounced, demonstrating superior memory efficiency for longer sequences.
-Additionally, Chronos-Offload further enhances ChronosPipe’s memory efficiency by reducing model state memory, allowing ChronosPipe to excel as both sequence length and model size scale up.
-In terms of throughput, while Chronos-Pipe experiences a relative throughput drop of 6%–9% compared to 1F1B, both Chronos-Recomp, and ChronosPipe ALL maintain throughput comparable to 1F1B+R=50% across all sequence lengths, even when throughput across sequence lengths are slightly influenced by Operator Library.
-Figure 12.
-estimated memory for different models at (PP,TP)=(8,8), global batch=128, micro batch=2.
-Ideal Computation Fraction is defined with following equation:(1- bubble overhead - recomputation overhead).
-As shown in Fig.
-12, we further evaluated the performance of ChronosPipe on dense models under multiple different parameter configurations.
-Chronos-Pipe+Chronos-Recomp achieves a storage reduction of 1.21x-1.26x while maintaining a throughput similar to that of 1F1B+R=50%.
-This enables the training of PaLM-62B and OPT-66B models within the 32GB HBM constraint.
-Moreover, ChronosPipe ALL achieves a storage reduction of 1.56x-1.58x compared to 1F1B+R=50%.
-Compared to 1F1B+R=100%, ChronosPipe not only boosts the throughput by approximately 1.15x but also reduces the storage by 1.04x-1.10x.
-6.3. Further Analysis of ChronosPipe
-Figure 13.
-analysis on P2P overhead on 24-layer model at (PP,TP)=(4,8), micro batch=2, seq len=4K.
-Ideal Computation Fraction is defined with following equation:(1- bubble overhead - recomputation overhead).
-As illustrated in Fig.
-13, we evaluated the impact of P2P communication on ChronosPipe under a PP4_TP8 configuration with a 24-layer model.
-During execution, Chronos-Pipe showed approximately a 6% lower ideal computation fraction compared to 1F1B.
-The theoretical analysis attributes 5% of this reduction to the overhead introduced by P2P communication, indicating that the primary gap stems from the additional round of P2P communication.
-This significant P2P overhead can be attributed to low P2P bandwidth and frequent P2P communication(one P2P communication happens after every three layers of computation in this setup), leading to $T_{c}\approx 0.104T_{unit}$.
-Despite this, Chronos-Recomp achieves an ideal computation fraction comparable to 1F1B+R=50% ($\leq 3\%$), suggesting that its overall P2P cost is similar to that of 1F1B.
-Specifically, Chronos-Recomp overlaps roughly half of the forward pass P2P communication with computation.
-However, due to the limited number of layers in the workload, gradient P2P communication during the backward pass (in chunk 2) does not overlap effectively with recomputation, resulting in minor degradation.
-In practice, ChronosPipe’s superior memory efficiency allows it to train larger workloads, naturally reducing the relative impact of P2P communication.
-As the workload scales, Chronos-Recomp achieves better overlap for gradient P2P communication during the backward pass, further mitigating its impact on performance.
-Figure 14. evaluate the scalability of Chronos-Offload on a 16-layer model at global batch=128, micro batch=2.
-We further verify the scalability of Chronos-Offload, as illustrated in Fig.
-14.
-When training a 16-layer model with a sequence length of 4K on PP4_TP8 setup, Chronos-Pipe achieves only partial overlap (45.45%) between the bubble created during the cooldown phase and the time required for weight gradient offloading and CPU-based optimizer updates.
-This limited overlap results in a slight increase in overall training time.
-However, when either the pipeline stage or the sequence length is doubled, the overlap improves significantly to 94.55% and 100%, respectively.
-Furthermore, Chronos-Offload enables only half of the optimizers to be updated on GPU, leading to a slight decrease in overall training time.
-These results highlight the scalability of Chronos-Offload, demonstrating its increasing effectiveness as the PP stage and sequence length grows, making it a viable solution for handling larger workloads.
-6.4. Further Design Space Exploration
-Figure 15. in-depth analysis of Chronos-Recomp. Additional memory required for recomputation is ignored for better illustration.
-Fig.
-15 illustrates our in-depth analysis of Chronos-Recomp.
-We theoretically evaluated Chronos-Recomp performance under PP4_TP8 configuration with chunk sizes of 2, 3, and 4.
-While small PP is less favorable for larger chunks in Chronos-Pipe, a clear trend is observed: Chronos-Recomp prioritizes recomputation tasks based on their temporal locality.
-When the recomputation budget is limited, focusing recomputation on shallower layers provides higher gain.
-For example, with a chunk size of 4, recomputing only 25% of the layers can reduce activation memory usage by up to 43.75%.
-However, as the recomputation ratio increases, Chronos-Recomp incurs a higher cost for recomputation in layers where it is less efficient.
-Nevertheless, Chronos-Recomp consistently outperforms standard recomputation strategies, as shown in Fig.
-15.
-Figure 16. in-depth analysis of Chronos-Offload.
-When partitioning the workload into more chunks, Chronos-Offload has the potential to offload more model states.
-Our theoretical analysis, illustrated in Fig.
-16, explores this behavior while maintaining throughput.
-The results indicate that Chronos-Offload experiences diminishing returns as the number of chunks increases; in some cases, memory efficiency can even decrease.
-This effect arises because adding more chunks only marginally increases the number of usable bubbles in the cooldown phase, particularly under smaller PP configurations.
-For example, when the PP stage is set to 4, using chunk=3 only provides 50% more bubbles compared to chunk =2, while chunk =4 produces the same number of bubbles as chunk =3.
-Consequently, the advantages of chunk =3 are mainly evident with long sequence length, whereas chunk =4 underperforms because the number of chunks matches the PP stage, ultimately reducing overall memory efficiency.
-7. RELATED WORK
-Hybrid Parallelism.
-To address the challenges of scaling model size and sequence length, modern distributed training systems leverage GPU clusters to expand overall memory capacity while deploying various schedules.
-These strategies include data parallelism(DP), pipeline parallelism(PP), tensor parallelism(TP), context parallelism(CP), and so on.
-DP (ZeRO, ; FSDP, ), PP (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ), and TP (AccPar, ; 2D-TP, ; 2p5D-TP, ; 3D-TP, ) enable the distribution of model states across machines.
-CP (RingAttention, ; Ulysses, ) and sequence parallelism(SP) (SeqParallel, ) specialize in distributing activations across their respective dimensions.
-Hybrid parallelism combines these approaches, enabling efficient scaling to models with trillions of parameters.
-PP is an essential component of hybrid parallelism among these strategies, particularly in cross-node deployments, due to its low communication requirement.
-However, PP suffers from pipeline bubbles and high activation memory overhead.
-Moreover, it is often less compatible with advanced DP like ZeRO-2/3.
-Pipeline Parallelism.
-Most pipeline parallelism research focuses on reducing pipeline bubbles by finer-grained task partitioning  (GPipe, ; DAPPLE, ; Interleave-1F1B, ; Hanayo, ; ZeroBubble, ) or multi-dimensional task scheduling  (Chimera, ; Mixpipe, ).
-However, they neglect the significant overhead of activation memory.
-(CtrlMemPP, ) reduces activation lifespans through scheduling but risks introducing more bubbles and prolonging gradient lifespan.
-It’s V-shaped scheduling also complicates compatibility with ZeRO-2/3 and Chronos-Offload, making storage optimization for other data more challenging.
-(AdaPipe, ) addresses activation memory imbalance with asymmetric partitioning but increases programming complexity.
-(BreadthFirstPP, ; ZeroPP, ), achieve ZeRO-2/3-compatibile PP at the cost of higher activation memory.
-(LLAMA3, ) takes advantage of  (BreadthFirstPP, ) and 1F1B for improved ZeRO-2 compatibility but only maintains, rather than reduces, activation memory usage.
-Recomputation and Offloading.
-In addition to optimizing multi-GPU systems’ HBM utilization through hybrid parallelism, memory requirements can also be reduced by trading memory for computation or communication time.
-For recomputation, some studies  (SeqParallel, ; MegatronKwai, ; SuperNeuron, ; CapuChin, ) focus on Selective recomputation of different operators but fail to address the high memory usage of the projection operator.
-(DTR, ) introduced temporal locality-aware recomputation principle and demonstrated its efficacy in DP, but coarse-grained PP and risk of inducing steady-phase bubbles limit the application of this principle in PP.
-Some approaches leverage the characteristics of Hybrid parallelism.
-(AdaPipe, ) adapts recomputation to mitigate activation memory imbalance in PP.
-(OverlapRecomp, ) overlaps recomputation with collective communication in the TP dimension to reduce overhead.
-However, these approaches increase programming complexity.
-For offloading,  (Bpipe, ) addresses memory imbalance in PP by offloading Activation across HBM.
-Most studies focus on fully utilizing storage systems.
-For example,  (ZeRO-Offload, ; PatrickStar, ; Mobius, ) explore offloading activation and model states to DRAM, while  (ZeRO-Infinity, ; Angel-ptm, ; Smart-Infinity, ; Fuyou, ) extend this to NVMe.
-Due to limited bandwidth in offloading, these efforts primarily focus on identifying opportunities to overlap offloading with computation and realize near-data processing.
-8. CONCLUSION
-This work introduces ChronosPipe to address the memory capacity limitations in LLM training.
-The core idea behind ChronosPipe is to treat HBM as a fast but limited ”cache,” optimizing and leveraging temporal locality in LLM pretraining for efficient HBM usage.
-Chronos-Pipe enhances activation temporal locality in PP scheduling and, together with Chronos-Recomp and Chronos-Offload, discards data with poor temporal locality from HBM.
-Experiments show that ChronosPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
-At the same time, ChronosPipe has good enough compatibility to make it suitable for large-scale hybrid parallelism pre-training.
-Appendix A Formalization on Delay Rounds at Chronos-Recomp
-In Chronos-Pipe scheduling, an additional interval time $T_{fwd\_interval}$ exists between forward chunks, derived from steps 1, 2, and 3 in Fig.
-5(a), resulting in $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$.
-In Chronos-Recomp’s shallow-layer full recomputation, the execution times for steps 1, 2, and 3 respectively change $T_{unit},\lceil\frac{(p-3)}{6}\rceil T_{unit}$, and $\lceil\frac{(p-1)}{2}\rceil T_{unit}$.
-Therefore, $\Delta T_{fwd\_interval}$ becomes $(\lceil\frac{(p-1)}{2}\rceil-\lceil\frac{(p-3)}{6}\rceil-1)T_{unit}(P\geq 3)$.
-Based on this, We can formalize the number of rounds k required to delay the launch of chunk 2 as an optimization problem for $P\geq 3$:
-(8) $Min$ $\quad k$
-$s.t.\quad T_{fwd\_interval}-$ $\Delta T_{fwd\_interval}+7kT_{unit}\geq 0$
-$k$ $\in N$
-$p$ $\geq 3$
-Therefore, when $8\leq p\leq 40$, we have $k=1$, meaning that chunk 2 needs to be delayed by one round.